### v1

# Lite-Mind: Towards Efficient and Versatile Brain Representation Network

Abstract
Research in decoding visual information from the brain, particularly through the non-invasive fMRI method, is rapidly progressing. The challenge arises from the limited data availability and the low signal-to-noise ratio of fMRI signals, leading to a low-precision task of fMRI-to-image retrieval. State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a deep MLP with a high parameter count orders of magnitude, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s vision transformer. However, significant individual variations exist among subjects, even within identical experimental setups, mandating the training of subject-specific models. The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices, especially with the necessitating of specific models for each subject. To this end, we propose Lite-Mind, a lightweight, efficient, and versatile brain representation network based on discrete Fourier transform, that efficiently aligns fMRI voxels to fine-grained information of CLIP. Our experiments demonstrate that Lite-Mind achieves an impressive 94.3% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye. Lite-Mind is also proven to be able to be migrated to smaller brain datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset. Code is available at https://github.com/gongzix/Lite-Mind.
1 Introduction
Brain decoding holds immense significance in elucidating intricate mental processes and serves as a pivotal technology for brain-computer interfaces [24, 14, 26]. Among diverse brain decoding tasks, the decoding of visual information stands out as a paramount yet challenging endeavor, allowing us to unravel the complex mechanisms involved in visual processing, object recognition, scene understanding, and even higher-order cognitive functions. In the pursuit of decoding natural visual information, functional magnetic resonance imaging (fMRI) is widely employed as a non-invasive modality to decipher the perceptual and semantic information intricately encoded within the cerebral cortex [20]. It draws considerable attention to fMRI-to-image retrieval/reconstruction tasks. This paper focuses primarily on tasks related to fMRI-to-image retrieval.
Figure 1: The significance of lightweight brain decoding model. Clouds of different colors represent different subject-wise models, and the size of the cloud represents the size of the model. The storage cost of MindEye is extremely high, and each person’s model parameter quantity is at the GB level, while Lite-Mind only requires the MB level parameter quantity.
The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images. However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models. To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models [27, 15, 9, 1, 13, 44]. Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models [28] as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures [21, 23, 19, 33]. Initially, the CLS token of the CLIP image encoder was used to align voxels in Mind-Reader [19]. Considering the CLS token alone provides limited information for fMRI data, BrainClip [21] and UniBrain [23] introduce the text annotations of images to assist in aligning fMRI voxels with image representations. Brain-diffuser [25] aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
Among the CLIP-related models, MindEye [33] stands out by employing a large-scale (996M) MLP Backbone and Projector to map fMRI voxels and align their representations with the final hidden layer of CLIP using contrastive learning. MindEye achieves remarkable breakthroughs, surpassing $90\%$ accuracy for both image and brain retrieval, outperforming previous state-of-the-art accuracies that were below $50\%$. However, the practical application of large-scale models in brain science research faces limitations. Each subject’s fMRI data can exhibit significant variations, even within identical experimental setups, making it currently infeasible to establish a universal encoding model that performs well for all individuals. Consequently, training individual encoding models for each subject becomes necessary, as depicted in Figure 1. Apparently, it is impractical to expect hospitals to have access to computing resources with sufficient performance to train a unique large-scale model (e.g., MindEye) for each individual. Therefore, the ideal scenario involves equipping individuals with lightweight brain-computer interface edge devices, enabling image representations to be efficiently retrieved by a lightweight model. To this end, there is an urgent need to develop a brain network that is lightweight and efficient, as illustrated in the right segment of Figure 1, for practical implementation and deployment of brain-computer interfaces in real-world scenarios.
MindEye strongly benefits from an MLP backbone with a high parameter count, which allows for straightforward and powerful compression and preservation. However, when dealing with fMRI data, which typically has a low signal-to-noise ratio, an MLP focusing on voxel value-wise mappings may inadvertently suppress informative voxel signal values while being inefficient at reducing noisy values. This can lead to model parameter redundancy and decelerated convergence speed. Recently, the Fourier transform has gained attention in deep learning, demonstrating lightweight characteristics and improved efficiency in learning frequency domain patterns, due to its global perspective and energy concentration properties [38]. Additionally, the Fourier transform naturally possesses the advantage of processing signals: it is more effective and efficient to filter out noise and maintain informative voxel signals in the frequency domain. It presents a promising alternative to large-scale MLPs for encoding fMRI signals.
To tackle the trade-off between model capacity and scalability in fMRI-to-image retrieval, we propose Lite-Mind, a lightweight, efficient, and versatile brain representation network. We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the heavy MLP Backbone used in MindEye. Extensive experiments show that Lite-Mind achieves 94.3% retrieval accuracy for Subject 1 on the NSD dataset, with 98.7% fewer parameters than MindEye. Lite-Mind also proves its adaptability to smaller brain datasets and establishes a new state-of-the-art zero-shot classification on the GOD dataset. Our contributions are summarized as follows:
  • We empirically show that it is unnecessary to overly focus on each voxel value of fMRI-to-image retrieval while focusing on the global features of fMRI patches is more lightweight and efficient.
  • We propose a new Brain Representation Network DFT Bcakbone based on Discrete Fourier Transform, as an efficient and versatile fMRI representation network for fMRI-to-image retrieval. Its effectiveness, efficiency, and lightweight are demonstrated.
  • We migrated DFT Backbone to a smaller GOD dataset and demonstrated its universality. It achieves state-of-the-art zero-shot classification performance.
2 Related Work
2.1 Brain Visual Decoding
The study of visual decoding using fMRI in the human brain has been a long-standing endeavor. In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation of performance is how to align fMRI signals to the image representation space or intermediate space. Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes. Further research utilizes pre-trained VGG to enrich hierarchical image features [13, 34, 7]. With the development of cross-modal tasks, the image-text pre-training model CLIP has been introduced into fMRI-image research. Mind-reader [19] is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow resnet-like model to align fMRI with the CLS token output by the CLIP image encoder. Mind-Reader did not achieve good performance in fMRI-image retrieval when using InfoNCE loss. BrainClip [21] uses a VAE model for retrieval tasks while still using the CLS token. Brain diffuser [25] introduced a fine-grained representation of the image of the last hidden layer in CLIP. But they used 257 different linear regression models to align each layer of the 257 hidden layers, resulting in the model being too cumbersome. Mind-Vis [5] conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values. Using a linear regression model, Takagi et al. [36] introduce additional ventral brain regions and align fMRI to the Stable Diffusion variable space. MindEye [33] uses a deep MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 [29] for the first time to narrow the disjointed vectors. At the same time, it creatively proposed a mixed contrast loss MixCo as a data augmentation method to make the large model converge. However, to the best of our knowledge, no previous research has focused on lightweight and efficient brain networks.
2.2 Fourier Transform in Deep Learning
Fourier transform plays a vital role in the area of digital signal processing. It has been introduced to deep learning for enhanced learning performance [11, 8, 43]. GFNet [30] utilizes fast Fourier transform to convert images to the frequency domain and exchange global information between learnable filters. As a continuous global convolution independent of input resolution, Guibas et al. [12] design the adaptive Fourier neural operator frame token mixing. Xu et al. [41] devise a learning-based frequency selection method to identify trivial frequency components and improve the accuracy of classifying images. On text classification, Lee-Thorp et al. [18] use the Fourier transform as a text token mixing mechanism. Furthermore, the Fourier transform is also applied to forecast time series  [3, 17, 16, 42]. To increase the accuracy of multivariate time-series forecasting, Cao et al. [3] propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the spectral domain. Yang et al. [42] propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum. Wang et al. [38] prove that frequency domain MLP is more efficient than real domain MLP. fMRI is essentially a signal that can be processed in the frequency domain, but to the best of our knowledge, there has been no research on applying FFT to efficient brain representation.
3 Preliminaries
3.1 Motivation and Analysis
FMRI signals are considered to have spatial redundancy, noise, and sample sparsity; even the nsdgeneral of NSD data has been preprocessed, which makes huge fully connected MLP of MindEye consume a lot of unnecessary parameter quantities. Furthermore, various data augmentation methods in MindEye are only used to prevent the overfitting of large models, which is also caused by MLP’s excessive focus on each voxel value. The fine-grained features come from the rich representation of the image by the last hidden layer of CLIP. Aligning multiple representations does not require excessive attention to each voxel’s values but rather to the different patterns of the entire brain region.
The diffusion model that approximates disjointed cross-modal vectors is effective for retrieval. This paper does not discuss the lightweight of the diffusion model but rather discusses the efficient brain representation network backbone. We focus on designing a lightweight voxel encoder that can quickly capture global information of fMRI and efficiently align fMRI voxels into fine-gained image representations.
4 Lite-Mind
Figure 2: Overview of our proposed Lite-Mind. Figure (a) is the architecture of the MLP Backbone of MindEye with 4 residual blocks(4096×4096). Figure (b) is our architecture of the DFT backbone and pipeline of retrieval. FMRI voxels are represented through DFT Backbone to obtain voxel embeddings. On the upper right is the overall retrieval pipeline, where LAION-5B retrieval and zero-shot classification are based on CLS token and diffusion projector. At the bottom right is the filter pyramid aggregation method, reducing the receptive field to align CLIP representation.
4.1 Overview
As illustrated in Figure 2, Lite-Mind comprises three components: 1) DFT Backbone maps flattened voxels to an intermediate space with a dimension of 257 × 768, corresponding to the last hidden layer of CLIP ViT/L-14. 2) Retrieval pipeline contains various downstream retrieval-based tasks with the voxel embeddings from the DFT backbone. 3) Contrastive Learning aims to bring the representation of two modalities closer together.
4.2 DFT Backbone
The backbone consists of patch partitioning followed by multilayer filter blocks, while the filter block adopts a pyramid aggregation strategy and finally outputs voxel embeddings. Given a fMRI-image sample pair [$x_{i}$,$y_{i}$] ($x_{i}$ represents the $i$-th fMRI sample, $y_{i}$ represents the image paired with it). Our optimization objectives are as follows:
[…]
Patchfy. Given a fMRI-image pair $x_{i}=[v_{1},v_{2},...,v_{m}]$ ($i$ represents the $i$-th fMRI sample, $m$ is the length of fMRI voxels). we firstly divide $x_{i}$ into $n$ non-overlapping region patches $p_{i}=[p_{1},p_{2},...,p_{n}](m=nl)$ and adopt CNN to generate the multi-channel representation of voxel patches:
$p^{c}_{i}=PE(CNN(x_{i}))$ (2)
Where $PE(\cdot)$ is the region patch embedding for voxels, and $c$ is the number of patch channels.
Filter Block. After we obtain the voxel patch embeddings $x_{i}=[p^{c}_{1},p^{c}_{2},...,p^{c}_{n}]$, then the spectrum of voxel patch embeddings is obtained by multi-channel 1D DFT as below:
$X^{c}[k]=F_{pat}(x[i])=\sum_{i=0}^{n-1}x[i]e^{-j(2\pi/m)ki}$ (3)
Where $X\in\mathbb{C}^{n\times d}$ is a complex tensor, $F_{pat}(\cdot)$ denotes the 1D DFT along the voxel patch dimension. Note that self-attention computes the spatial dependencies in a quadratic time complexity, while DFT can be efficiently implemented via a fast Fourier transform in logarithmic time complexity.
After obtaining the spectrum of voxel patches, we use a voxel filter library for filtering processing. Voxel spatial features are effectively consolidated within each frequency element, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain. To expand voxel vectors to final hidden layers of CLIP with dimensions much larger, we introduce a filter library for each channel of voxel patches $X^{c}$, to obtain fine-grained alignment of images in the frequency domain. We use $K=[K_{c_{1}},K_{c_{2}},...,K_{c}]$ to represent the filter library, where c is the number of filters in the filter library:
$\hat{X}=\sum_{i=1}^{c}\frac{1}{l}X^{i^{2}}\odot k_{i}cos(\frac{(2i-1)\pi}{2k})$ (4)
where $\odot$ is the element-wise multiplication, $|X^{i}|^{2}$ is the power spectrum of $X^{i}$, $l$ is the length of $X^{i}$. The $|X^{i}|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective. It also facilitates voxel information concentration and noise filtering. $cos((2i-1)\pi/2k)$ compacts better energy and can aggregate the more important information in fMRI voxels. Its combination of application with the channel filter library $K$ allows for various fine-grained features alignment to images from different channel modes.
Filter Pyramid Merge. To better align the voxel patches with the fine-grained representation of the image in the frequency domain, we adopt a filter pyramid merging strategy to expand the receptive field of the voxel patch channels. Gradually attenuated filter library converges to obtain final hidden layer $257\times 768$ for alignment, as shown in the lower right corner of Figure 2.
4.3 Retrieval pipeline
The retrieval process on various downstream tasks is shown in the upper right of Figure 2.
Test set retrieval. The cosine similarities between voxel embeddings and image embeddings are directly calculated.
LAION-5B retrieval. Image embeddings are generated based on voxel CLS token by Diffusion Projector for online retrieval. Because our DFT backbone uses the same one-way contrast loss as CLIP, we consider that the DFT backbone outputs voxel embeddings similar to text embeddings, which we consider as voxel-level text. We found that when searching on the large-scale dataset LAION-5B, img2img performs better than text2img and can find more similar images. Therefore, we used a similar method to represent the disjointed voxels of the backbone output, using a DALLE · 2[29] similar diffusion model to obtain the image representation and retrieve it on LAION-5B.
Zero-shot classification. Image retrieval is performed on novel classes in the test set, and the similarities between the retrieved images and simple CLIP class text prompts are calculated for classification tasks.
4.4 Contrastive Learning
Due to the inherent differences between different modalities, we use CLIP’s contrastive loss to train the DFT backbone. To perform img2img retrieval on LAION-5B, we use MSE loss to constrain the generation of approximate image representations. Sampling a batch $\mathcal{B}$ from voxel-image pairs, the contrastive loss is defined below:
$L_{contr}=-\frac{1}{|\mathcal{B}|}\sum_{s=1}^{|\mathcal{B}|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|\mathcal{B}|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$ (5)
where $f_{s}$ is the $s$-th voxel representation, $\mathcal{V}_{s}$ is its corresponding image representation, $\tau$ is a temperature factor.
Accordingly, the training objective is defined below:
$L=\alpha L_{contr}+\beta L_{mse}$ (6)
When training LAION-5B retrieval tasks, $\beta$ is a non-zero value; When training test set retrieval, $\beta=0$. Due to the lightweight performance of the model, we trained two separate models for fMRI-image retrieval and image-fMRI retrieval, corresponding to two types of one-way losses.
5 Experiments
5.1 Dataset
Natural Scenes Dataset (NSD) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images from MSCOCO-2017 dataset [1], which contains images of complex natural scenes. Participants viewed three repetitions of 10,000 images with a 7-Tesla fMRI scanner over 30–40 sessions. More details can be found on the NSD official website [^1] https://naturalscenesdataset.org. Our experiments focused on Subj01, Subj02, Subj05, and Subj07, who finished all viewing trials. Each subject’s training set comprises 8859 image stimuli and 24980 fMRI trials (with the possibility of 1-3 repetitions per image). The test set contains 982 image stimuli and 2770 fMRI trials. Responses for images with multiple fMRI trials are averaged across these trials. By applying the nsdgeneral ROI mask with a 1.8 mm resolution, we obtained ROIs for the four subjects, comprising 15724, 14278, 13039, and 12682 voxels respectively. These regions span visual areas ranging from the early visual cortex to higher visual areas. Our experimental setup is consistent with the NSD image reconstruction and retrieval articles [36, 25, 23, 33, 19].
Generic Object Decoding (GOD) Dataset [^2] http://brainliner.jp/data/brainliner/Generic_Object_Decoding was created by Horikawa and Kamitani [13], consisting of fMRI recordings of five healthy subjects who were presented with images from ImageNet [6]. The GOD Dataset includes 1250 distinct images selected from 200 ImageNet categories. Among these, 1200 training images are drawn from 150 categories, and 50 test images are from the remaining 50 categories. The training and test image stimuli were presented to the subjects once and 35 times respectively, resulting in 1200 and 1750 fMRI instances. We used preprocessed ROIs, encompassing voxels from the early visual cortex to higher visual areas. For each test image, the fMRI responses from different trials were averaged.
Figure 3: Partial search results of Lite-Mind on all 982 test images for Subject 1. Even without diffusion prior and heavy MLP backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates. The number below each image represents the similarity score.
Method Model Parameters Retrieval
[…]
Lin et al. [19] deep models 2.34M 11.0% 49.0%
Ozcelik… [25] 257 separate linear regression models 3B 21.1% 30.3%
MindEye [33] MLP Backbone 940M 13.3% 63.1%
MindEye [33] MLP Backbone+Projector 996M 88.8% 84.9%
MindEye [33] MLP Backbone+Prior 3B 93.4% 90.1%
Lite-Mind(ours) DFT Backbone 12.5M 94.3% 97.1%
(a) Quantitative comparison of Lite-Mind retrieval performance against other models for Subject 1. Image retrieval refers to the percent of the time the correct image was retrieved out of 300 candidates, given the associated brain sample (chance=0.3%); vice-versa for brain retrieval. Lite-Mind only uses a parameter quantity of 12.5M, achieving extremely high retrieval performance without using any model to close vectors of different modalities(see Appendix B for remaining subject models).
Method Low-Level High-Level
[…]
Lin et al. [19] - - - - 78.2% - - -
Takagi… [36] - - 83.0% 83.0% 76.0% 77.0% - -
Gu et al. [10] .150 .325 - - - - .862 .465
Ozcelik… [25] .254 .356 94.2% 96.2% 87.2% 91.5% .775 .423
BrainCLIP [21] - - - - 86.7% 94.8% - -
MindEye [33] .309 .323 94.7% 97.8% 93.8% 94.1 .645 .367
MindEye-125M(LAION-5B) .130 .308 84.0% 92.6% 86.9% 86.1% .778 .477
Lite-Mind-8M(LAION-5B) .125 .331 78.7% 89.4% 87.9% 88.7% .724 .446
(b) LAION-5B retrieval performance. The upper part is based on various fMRI-image reconstruction methods using generative neural networks, while the lower part represents the retrieval performance of the large-scale dataset LAION-5B as an alternative to fMRI-image reconstruction. All results are averaged across four subjects, and Lite-Mind achieved higher performance with a lighter backbone, even exceeding the accuracy of many reconstruction methods.
5.2 Implementation details
To fairly compare different backbones, we did not use any data augmentation methods, such as voxel mixing loss MixCo and image slicing enhancement. Due to MindEye only disclosing the MLP Backbone performance of Subject 1, only the performance of Subject 1 is provided in Table 0(a). The experimental results of more subjects are shown in Appendix B. All of our experiments were conducted on a single Tesla V100 32GB GPU. More experimental details and hyperparameter settings can be found in Appendix A.
6 Results
6.1 fMRI/image retrieval
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings in the training set. If a paired image embedding is retrieved, the retrieval is considered correct. fMRI retrieval is the opposite process mentioned above. Note that there are many semantically and visually similar images in the NSD dataset test set, which are considered to be similar in CLIP space. Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image. For test retrieval, we adhered to the identical methodology as Lin et al. [19] and MindEye [33] to compute the retrieval metrics presented in Table 0(a). For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images. The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 982 voxel embeddings of the test set. We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples. The experimental results are shown in Table 0(a). We also conducted retrieval experiments on the remaining three subjects to demonstrate the universality of DFT Backbone. The detailed results and discussions are in Appendix B.
As shown in Table 0(a), compared to MLP Backbone+Projector our DFT Backbone improves retrieval accuracy by 5.5% and 12.2% for two retrieval ways, indicating that the fine-grained representation of the image comes from the rich representation of the last hidden layer of CLIP rather than MLP’s excessive attention to each voxel value. Furthermore, our DFT Backbone retrieval performance even surpasses MindEye’s MLP Backbone+Prior without any image enhancement methods, proving that frequency domain calculations better capture global features.
Figure 4: Partial search results of Lite-Mind on LAION-5B retrieval for Subject 1. The left column marked by a red box in every two columns represents the original image seen by the subject, and the right column represents the image retrieved by LAION-5B.
Methods Modality Prompt Subject1 Subject2 Subject3 Subject4 Subject5
top-1 top5 top-1 top5 top-1 top5 top-1 top5 top-1 top5
CADA-VAE [31] V&T - 6.31 35.70 6.45 40.12 17.74 54.34 12.17 36.64 7.45 35.04
MVAE [40] V&T - 5.77 31.51 5.40 38.46 17.11 52.46 14.02 40.90 7.89 34.63
MMVAE [35] V&T - 6.63 38.74 6.60 41.03 22.11 56.28 14.54 42.45 8.53 38.14
MoPoE-VAE [22] V&T - 8.54 44.05 8.34 48.11 22.68 61.82 14.57 58.51 10.45 46.40
BraVL [7] V&T - 9.11 46.80 8.91 48.86 24.00 62.06 15.08 60.00 12.86 47.94
BrainClip-VAE [21] V&T Text 8.00 42.00 24.00 52.00 20.00 58.00 20.00 58.00 20.00 46.00
BrainClip-VAE [21] V&T CoOp 14.00 54.00 20.00 60.00 21.33 64.67 16.67 66.67 18.00 52.00
MindEye-211M [33] V Text 6.00 40.00 20.00 60.00 28.00 70.00 14.00 54.00 10.00 64.00
Lite-Mind-15.5M(ours) V Text 26.00 74.00 28.00 70.00 30.00 80.00 34.00 76.00 24.00 66.00
Table 1: Zero-shot visual stimulus classification on GOD dataset. The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%). The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from [7]. MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B for retrieval performance on the GOD dataset).
6.2 LAION-5B retrieval
Because Lite-Mind can also align image representation space, we can expand the retrieval task to larger datasets, such as LAION-5B [32]. For LIAON-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS token of CLIP. The final layer CLIP ViT-L/14 embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client [2]. For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS token, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP). Unlike MindEye, in the LAION-5B retrieval process, we used the diffusion model of DALLE·2 prior to conducting img2img retrieval.
To compare the ability of LAION-5B retrieval to replace image reconstruction, we compared the performance indicators of reconstruction with previous image reconstruction methods. The experimental results are shown in Table 0(b). The specific evaluation indicators in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstructions; SSIM is structural similarity index metric [39]; EfficientNetB1(”Eff”) [37] and SwAV-ResNet50(”SwAV”) [4] refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
Experiments show that our diffusion projector performs better than the projector of MindEye in high-level metrics, as shown in Table 0(b), although our backbone only has a parameter size of 8M (MindEye-CLS backbone still has 125M). The results indicate that efficient DFT backbones can also assist downstream retrieval tasks. Partial search results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval performance and results of each subject are shown in Appendix B.
6.3 GOD zero-shot classification
Since the collective prime values and data volume of fMRI data are much smaller than those of the NSD dataset in most cases, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify the Lite-Mind’s generalization ability. The voxel values collected by participants in the GOD dataset range from 4133 to 4643, much smaller than that of the NSD dataset, and the training set has only 1200 fMRI-image pairs(4.8% of NSD training set).
We conduct zero-shot classification by using fMRI-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification). We train our Lite-Mind-15.5M to align voxels to images on the GOD dataset. We also replicated the MLP backbone and projector methods of MindEye, changing the residual block of $4096\times 4096$ to $1024\times 1024$ to accommodate the reduction in voxel vector length, and MindEye still has a parameter quantity of 210M. As shown in Table 1, the experimental results indicate that our proposed Lite-Mind still guarantees versatility and establishes a new state-of-the-art for zero-shot classification on the GOD dataset without complex text prompt templates or auxiliary text representation training. On the contrary, due to the sudden decrease in data size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data volume and data augmentation for voxel value-wise mappings in heavy MLP Backbone. More retrieval and classification results are provided in Appendix B.
6.4 Ablations and visualization
Architectural Analysis. We assess the effectiveness of different modules within Lite-Mind to investigate the lightweight and efficient performance in this session. All ablation experiments are for Subject 1 on the NSD dataset.
We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of DFT on retrieval efficiency(Table 2). We even completely removed the DFT layer and aligned the patches directly to the last hidden layer of CLIP ViT-L/14 with a tiny MLP, still achieving a forward retrieval accuracy of 65.8%, indicating that patches have more concentrated information than voxels. As the number of DFT layers gradually deepens, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual. We interestingly found that the accuracy of brain retrieval tends to converge faster than the image and only 6 layers of DFT are needed to approach the highest value. We have also conducted more ablation experiments on data volume to demonstrate the robustness of Lite-Mind. Refer to Appendix B for more ablations.
Visualization. We visualize the diffusion model representation of LAION-5B using T-SNE in Figure 5. The graph shows that the diffusion model successfully transformed the voxel representation learned from contrastive learning into an image representation, and the two representations were well fused together for img2img retrieval on the LAION-5B dataset. We have also conducted more visualization experiments in Appendix C to demonstrate the efficiency of DFT Backbone.
DFT depth Parameters Image Retrieval Brain Retrieval
Pyramid(18) 12.5M 94.3% 97.4%
18 9.9M 91.8% 97.3%
12 6.6M 91.0% 97.2%
6 3.4M 90.6% 97.2%
1 0.7M 83.5% 94.0%
0 0.2M 65.8% 74.7%
Table 2: Ablation experiments on Discrete Fourier Transform (DFT) layers revealed a positive correlation between the depth of the DFT layer and enhanced forward retrieval accuracy. Simultaneously, augmenting the filter pyramid of the receptive field demonstrated an additional improvement in retrieval accuracy, provided there is a judicious increase in the parameter count.
Figure 5: T-SNE of representations of 982 test fMRI voxels by diffusion projector on LAION-5B retrieval for Subject 5. The diffusion model plays a role in bringing vectors closer, enabling img2img retrieval on the LAION-5B dataset.
7 Conclusion
In this work, we propose Lite-Mind, an extremely lightweight brain representation extraction network based on Fourier transform for cross-modal fMRI-image retrieval. Our DFT Backbone is an efficient means of obtaining fine-grained representation alignment between fMRI signals and visual stimuli. With high retrieval accuracy, Lite-Mind can transform many downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification. Meanwhile, Lite-Mind’s dependence on data volume and voxel values is less than that of larger models, demonstrating excellent universality across different fMRI datasets.
For future work, we should explore the relationship between different sizes of patches and brain area structures and explore the impact of cortices other than the visual cortex, such as ventral, on retrieval accuracy. In addition, versatile cross-subject models should be studied to efficiently project fMRI representations of different subjects into the same joint space or transfer from subject models with larger sample sizes to models with sparse sample sizes. At the same time, how to use an efficient model to collaboratively train multiple fMRI datasets (across subjects) is also a promising direction that can be explored in the future.
Supplementary Material
Appendix A Additional Method Details
A.1 Implementation details
Data preprocessing. We downloaded the NSD dataset from the official website and used Takagi’s code to extract nsdgenal fMRI voxels, while Takagi extracted the stream region of the NSD dataset. We noticed that MindEye scaled the fMRI voxel values in advance, while we did not. The difference in fMRI voxels input data is shown in the example in Figure 6. NSD image files come from nsd-stimuli.hdf5 file and have a unified size of $425\times 425$. We did not perform any data augmentation on the image and straightly extracted the hidden layer representation (size of $257\times 768$) of the image through CLIP ViT-L/14 for training.
Hyper-parameters. On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and $\beta=0$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval. Owning to the lightweight nature of Lite-Mind, the batch size is set to 1500, and the learning rate is 1.16e-3. For Subject 1, 2, 5, 7, Pyramid is set to [512, 256, 128, 257] respectively, and the corresponding number of layers and the patch size are set to [2, 10, 2, 4] and [480, 440, 460, 450], respectively.
About LAION-5B retrieval, for DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, while the weight decay is 6.02e-2 for diffusion projector. CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, and $\beta=0.5$, the batch size is set to 80 and the learning rate is 1.16e-3 for DFT Backbone while 1.1e-4 for diffusion projector.
On the GOD dataset, the hyperparameters are the same as those on the NSD dataset, except for the batch size and patch size which are set to 1200 and 8 for all 5 subjects on the GOD dataset, respectively.
(a) MindEye
(b) Lite-Mind
Figure 6: The fMRI averaged activities of MindEye and Lite-Mind responding to the same image, respectively. The figures only visualize the activities of the first 1000 voxels for illustration.
Appendix B Additional Experiment Results
B.1 Additional Results of Retrieval
To demonstrate the applicability of our method, we also conducted experiments on three other subjects, i.e., Subject 2, Subject 5, and Subject 7, on the NSD dataset, and the experimental results are shown in Table 3. In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths, but instead used smaller patches. It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90, and the retrieval accuracies of Subjects 5 and 7 are also above 75, The results prove that DFT Backbone can efficiently work on different subjects. Note that MindEye only includes the results of the overall model in Subjects 2, 5, and 7 in its appendix, and it does not evaluate the effect of individual MLP Backbone. As a result, Table 3 does not present any evaluation results for MindEye.
Method Voxel Length Parameters Retrieval
[…]
Lite-Mind(Subj 1) 15724 12.51M 94.3% 97.4%
Lite-Mind(Subj 2) 14278 12.49M 93.4% 98.2%
Lite-Mind(Subj 5) 13039 12.47M 80.5% 86.3%
Lite-Mind(Subj 7) 12682 12.47M 75.0% 82.3%
Table 3: Addtional retrieval performance for individual subjects on 982 test images of the NSD dataset.
Method Low-Level High-Level
[…]
Lite-Mind(Subject 1) .134 .332 78.8% 88.9% 88.5% 88.8% .730 .451
Lite-Mind(Subject 2) .120 .328 78.0% 89.4% 86.3% 87.4% .730 .446
Lite-Mind(Subject 5) .123 .332 79.4% 90.0% 88.8% 89.9% .712 .440
Lite-Mind(Subject 7) .121 .331 78.7% 88.8% 87.8% 88.5% .723 .448
Table 4: LAION-5B retrieval alternative reconstruction performance for the specific subject.
B.2 Additional Results of LAION-5B Retrieval
In order to better reflect the retrieval performance of Lite-Mind on LAION-5B, we presented the performance indicators of LAION-5B retrieval substitution reconstruction for other subjects, i.e., Subject 2, Subject 5, and Subject 7, in Table 4, corresponding to the average performance of subjects in Table 0(b). Similarly, the visualization results of the other subjects in Figure 7 correspond to the image samples of Subject 1 in Figure 4. Based on the comprehensive table and graph, it can be found that Lite-Mind has good generalization on all four subjects, verifying the LAION-5B retrieval ability of Lite-Mind on different subjects. Meanwhile, as shown in Figure 7, the retrieval performance of LAION-5B completely depends on the retrieval accuracy of the CLS model in the test set, such that images retrieved incorrectly in the test set may also have retrieval bias on LAION-5B, for example treating a teddy bear as an image of a cat or dog as shown in Figure 7. However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS token models. In the future, it would be beneficial to explore models that improve the alignment of CLS tokens or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
Figure 7: Addtional retrieval results corresponding to Figure 4. The left column marked by a red box in every two columns represents the original image seen by the subject, and the right column represents the image retrieved on LAION-5B.
B.3 Additional Results of Zero-shot Classification
In the main body of the paper, we only demonstrated the zero-shot classification effect of Lite-Mind on the GOD dataset. The corresponding retrieval results of each Subject are shown in the Table 5 below.
[…]
top1 top5
Lite-Mind(Subj 1) 4466 15.50M 30.0% 60.0%
Lite-Mind(Subj 2) 4404 15.46M 38.0% 58.0%
Lite-Mind(Subj 3) 4643 15.64M 38.0% 72.0%
Lite-Mind(Subj 4) 4133 15.24M 42.0% 62.0%
Lite-Mind(Subj 5) 4370 15.43M 26.0% 54.0%
Table 5: Retrieval performance on the GOD dataset.
B.4 Additional Ablations
In this section, we delved into the dependence of Lite-Mind on data volume and feature length on the NSD dataset and explored the specific impact of different cerebral cortexes on retrieval accuracy, not just nsdgenal. All the experimental results on the NSD dataset are derived from Subject 1, with a retrieving pool size of 300.
B.4.1 Retrieval with varying dataset size
In order to better investigate the versatile representation of DFT Backbone, we conducted data volume experiments on the NSD dataset, as shown in Table 6. Among them, All data corresponds to all trials of every single Subject on the NSD dataset, and 24890 fMRI-image pairs are used for training; Half data refers to randomly removing half of the training data; 2-session refers to using only 1500 randomly selected fMRI-image pairs for training, which corresponds to the two scan sessions of the fMRI experiment. In the table, we report the results of MLP Backbone and diffusion models for MindEye. The experimental results indicate that even in the same NSD dataset, the dependence of Lite-Mind on the number of training samples is smaller than that of MindEye. This also proves that the lightweight model has a wider applicability, i.e., versatility.
Method Parameters Retrieval
[…]
All Data(MindEye) 3B 97.2% 94.7%
Half Data(MindEye) 3B 77.5% 60.8%
2-Sessions(MindEye) 3B 17.9% 12.0%
All Data(Lite-Mind) 12.5M 94.3% 97.4%
Half Data(Lite-Mind) 12.5M 83.4% 91.9%
2-Sessions(Lite-Mind) 12.5M 11.9% 14.1%
Table 6: Retrieval performance with varying dataset size for Subject 1 on the NSD dataset.
Figure 8: Lite-Mind’s training loss curve and testing accuracy curve for Suject 1, 2, 5, 7 on the NSD dataset. The testing accuracy is calculated based on a retrieval pool of size 982.
B.4.2 Retrieval with varying representation sizes
We conducted additional experiments to verify the impact of representation dimensions on retrieval accuracy, as shown in Table 7. Among them, two different CLIP models (i.e., CLIP ViT/B-32 and CLIP ViT/L-14) were used to extract image representation, with a total of four representation dimensions, by CLS token and the last hidden layer, respectively. From the results, it can be seen that the retrieval accuracy of Lite-Mind is higher when the representation dimension of the image is longer and the representation of the image is richer. The results verify that the fine-grained alignment we mentioned in the main text comes from the rich representation of the image, rather than the fMRI voxel value fully connected heavy MLP Backbone.
Method Representation Parameters Retrieval
[…]
ViT-B/32 CLS 512 6.7M 57.5% 61.8%
ViT-L/14 CLS 768 6.7M 60.3% 64.2%
ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
ViT-L/14 Hidden 257×768 12.5M 94.3% 97.4%
Table 7: Retrieval performance with different CLIP representation for Subject 1 on the NSD dataset.
B.4.3 Cerebral cortex
In order to explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s method to extract the stream region of the NSD dataset, covering the visual cortex of nsdgenal, and separately trained the retrieval model for each cortex (early, lateral, parietal, ventral). As shown in Table 8, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy, although with only 5917 voxels Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research. They believe that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving that fully connected backbones are unnecessary. Interestingly, for other cortical regions, both lateral and ventral have a good impact on retrieval accuracy, while parietal shows little impact. The findings indicate the necessity of further research on whether the accuracy improvement is related to the longer voxel length of these two regions.
Region Voxel Length Parameters Retrieval
[…]
early 5917 12.4M 85.0% 93.4%
lateral 7799 12.4M 26.9% 29.0%
parietal 3548 12.3M 9.6% 11.2%
ventral 7604 12.3M 18.1% 22.4%
Table 8: Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
Figure 9: Lite-Mind’s retrieval heatmap on the NSD dataset for Subject 1. The larger figure on the right represents 982 × 982’s heatmap, and the smaller figure on the left represents the 50 × 50 subgraph.
Appendix C Additional Visualization
C.1 Training Curve
We visualized the training process of all four subjects on the NSD dataset in Figure 8. As the training epochs increased, the loss of the training set rapidly decreased, while the accuracy of the test set rapidly increased and then showed a slow upward trend. Accuracy refers to the hit rate of correct retrieval from 982 test set images.
C.2 Retrieval Heatmap
We visualized the retrieval heatmap for Subject 1 on all 982 test images of the NSD dataset in Figure 9. It can be observed that the similarity is highest on the diagonal, and the color of the retrieved heat map is darker. It shows that Lite-Mind has effectively retrieved corresponding images, even if there are many similar images in the test set, which verifies the fine-grained ability of Lite-Mind.
C.3 Visualization in Frequency Domain
We visualized the weights of different patches in different layers of filters, as shown in Figure 10 and 11. Visualization is divided into the real part (left) and imaginary part (right) of filter weights. It can be observed that different channels have varying degrees of attention to different patches, and the frequency domain better captures this characteristic. Interestingly, the weight of the imaginary part for the first and last patches is almost always 0, indicating that the noise is distributed in these two patches.
(a) The real part and imaginary of patch weights for CLS token channel.
(b) The real part and imaginary of patch weights for 140th channel.
Figure 10: Weights visualization for DFT Backbone of $257\times 768$ representation length.
(a) The real part and imaginary of patch weights for the 1st channel.
(b) The real part and imaginary of patch weights for the 60th channel.
Figure 11: Weights visualization for DFT Backbone of 768 representation length.
C.4 Information Alignment
We visualized the T-SNE plot between the voxel embeddings output by DFT Backbone and the image embeddings of frozen CLIP as the accuracy of the test set improved, as shown in Figure 12. We can observe that as the training progresses, the retrieval accuracy of the test set improves, and the shape of voxel embeddings tends to be closer to image embeddings, indicating the success of contrastive learning.
Appendix D Efficiency Discussion
D.1 Theoretical Analysis.
Theorem. Suppose that H is the representation of raw fMRI voxel patches and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of a voxel patch series in the spatial domain is equal to the energy of its representation in the frequency domain. Formally, we can express this with the above notations:
$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$ (7)
Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the spatial/channel dimension, $f$ is the frequency dimension.
Proof.
Given the representation of raw voxel patch series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in either the $C$ dimension (channel dimension) or the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{\textbf{H}(v)}{\textbf{H}^{*}(v)}\,\mathrm{d}v$ (8)
where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v$ $=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v$ (9)
=*$=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f$
=*$=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f$
$=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$
Proved. ∎
Therefore, the energy of a voxel patch series in the spatial domain is equal to the energy of its representation in the frequency domain.
Complexity Analysis. For a fMRI voxel with a length of $m$, we divide it into $n$ $l$-length patches ($m=nl$) with $C$ channel numbers. Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP and DFT respectively, the middle layer dimension of MLP is $D_{1}$, and the alignment representation dimension is $N\times D_{2}$, where $N$ is the number of the last hidden layer of CLIP. The time complexity of MLP Backbone is $O(mD_{1}+L_{1}D_{1}^{2}+ND_{1}D_{2})$. For DFT Backobone, the time complexity of patchfy is $O(mC)$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nC\operatorname{log}n+nC)$. The time complexity of each layer’s tiny MLP is $O(8nC^{2})$. Thus the time complexity of the entire DFT Backbone is:
$O(mC+L_{2}(2nC\operatorname{log}n+nC+8nC^{2})+nD_{2})$ (10)
$=O(mC+nC\operatorname{log}nL_{2}+nC^{2}L_{2}+nD_{2})$
Quantitative analysis algorithm complexity for DFT Backbone is shown in Table 9.
Model FLOPs(G) Parameters Image Retrieval Brain Retrieval
MindEye 5.66 996M 88.8% 84.9%
Pyramid-18 0.80 12.5M 94.3% 97.4%
DFT-18 0.63 9.9M 91.8% 97.3%
DFT-12 0.43 6.6M 91.0% 97.2%
DFT-6 0.22 3.4M 90.6% 97.2%
DFT-1 0.05 0.7M 83.5% 94.0%
DFT-0 0.02 0.2M 65.8% 74.7%
Table 9: Quantitative complexity analysis for DFT Backbone with different layers and MLP Backbone+Projector for MindEye.
Figure 12: T-SNE visualization between the voxel embeddings output by DFT Backbone and the image embedding of frozen CLIP. Accuracy in the title refers to the hit rate of correct retrieval from 982 test set images and the blue dots represent voxel embeddings.

### v2

@@
-Lite-Mind: Towards Efficient and Versatile Brain Representation Network
+Lite-Mind: Towards Efficient and Robust Brain Representation Learning
 Abstract
@@
 The challenge arises from the limited data availability and the low signal-to-noise ratio of fMRI signals, leading to a low-precision task of fMRI-to-image retrieval.
-State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a deep MLP with a high parameter count orders of magnitude, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s vision transformer.
-However, significant individual variations exist among subjects, even within identical experimental setups, mandating the training of subject-specific models.
-The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices, especially with the necessitating of specific models for each subject.
-To this end, we propose Lite-Mind, a lightweight, efficient, and versatile brain representation network based on discrete Fourier transform, that efficiently aligns fMRI voxels to fine-grained information of CLIP.
-Our experiments demonstrate that Lite-Mind achieves an impressive 94.3% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye.
-Lite-Mind is also proven to be able to be migrated to smaller brain datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset.
-Code is available at https://github.com/gongzix/Lite-Mind.
+State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a large model, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s Vision Transformer(ViT).
+However, significant individual variations exist among subjects, even within identical experimental setups, mandating the training of large subject-specific models.
+The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices.
+To this end, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm based on Discrete Fourier Transform, that efficiently aligns fMRI voxels to fine-grained information of CLIP.
+Our experiments demonstrate that Lite-Mind achieves an impressive 94.6% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye.
+Lite-Mind is also proven to be able to be migrated to smaller fMRI datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset.
+Code will be available at https://github.com/gongzix/Lite-Mind.
 1 Introduction
@@
 This paper focuses primarily on tasks related to fMRI-to-image retrieval.
-Figure 1: The significance of lightweight brain decoding model.
-Clouds of different colors represent different subject-wise models, and the size of the cloud represents the size of the model.
-The storage cost of MindEye is extremely high, and each person’s model parameter quantity is at the GB level, while Lite-Mind only requires the MB level parameter quantity.
+Figure 1: The significance of lightweight brain decoding model. Subject-specific models are visually represented by clouds of distinct colors, and the cloud size correlates with model parameters.
 The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images.
@@
 Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models [30] as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures [23, 25, 21, 35].
-Initially, the CLS token of the CLIP image encoder was used to align voxels in Mind-Reader [19].
-Considering the CLS token alone provides limited information for fMRI data, BrainClip [21] and UniBrain [23] introduce the text annotations of images to assist in aligning fMRI voxels with image representations.
+Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader [21].
+Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip [23] and UniBrain [25] introduce the text annotations of images to assist in aligning fMRI voxels with image representations.
 Brain-diffuser [27] aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@
 It presents a promising alternative to large-scale MLPs for encoding fMRI signals.
-To tackle the trade-off between model capacity and scalability in fMRI-to-image retrieval, we propose Lite-Mind, a lightweight, efficient, and versatile brain representation network.
-We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the heavy MLP Backbone used in MindEye.
-Extensive experiments show that Lite-Mind achieves 94.3% retrieval accuracy for Subject 1 on the NSD dataset, with 98.7% fewer parameters than MindEye.
+To tackle the trade-off between model capacity and scalability in fMRI-to-image retrieval, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm.
+We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the large MLP Backbone used in MindEye.
+Extensive experiments show that Lite-Mind achieves 94.6% retrieval accuracy for Subject 1 on the NSD dataset, with 98.7% fewer parameters than MindEye.
 Lite-Mind also proves its adaptability to smaller brain datasets and establishes a new state-of-the-art zero-shot classification on the GOD dataset.
 Our contributions are summarized as follows:
-  • We empirically show that it is unnecessary to overly focus on each voxel value of fMRI-to-image retrieval while focusing on the global features of fMRI patches is more lightweight and efficient.
-  • We propose a new Brain Representation Network DFT Bcakbone based on Discrete Fourier Transform, as an efficient and versatile fMRI representation network for fMRI-to-image retrieval.
+  • We demonstrated that the Fourier Transform has the advantage of lightweight and efficient for fMRI representation learning in frequency domain.
+  • We proposed a new Brain Representation Network DFT Backbone based on Discrete Fourier Transform, as an efficient and robust fMRI representation network for fMRI-to-image retrieval.
 Its effectiveness, efficiency, and lightweight are demonstrated.
-  • We migrated DFT Backbone to a smaller GOD dataset and demonstrated its universality. It achieves state-of-the-art zero-shot classification performance.
+  • We migrated Lite-Mind to various downstream tasks, and demonstrated its robustness. It achieves state-of-the-art zero-shot classification performance on the GOD dataset.
 2 Related Work
@@
 The study of visual decoding using fMRI in the human brain has been a long-standing endeavor.
-In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation of performance is how to align fMRI signals to the image representation space or intermediate space.
-Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes.
+In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation is how to align fMRI signals to the image representation space or intermediate space.
+Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes [38].
 Further research utilizes pre-trained VGG to enrich hierarchical image features [13, 36, 7].
 With the development of cross-modal tasks, the image-text pre-training model CLIP has been introduced into fMRI-image research.
-Mind-reader [19] is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow resnet-like model to align fMRI with the CLS token output by the CLIP image encoder.
+Mind-reader [21] is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow resnet-like model to align fMRI with the CLS embeddings output by the CLIP image encoder.
 Mind-Reader did not achieve good performance in fMRI-image retrieval when using InfoNCE loss.
-BrainClip [21] uses a VAE model for retrieval tasks while still using the CLS token.
+BrainClip [23] uses a VAE model for retrieval tasks while still using the CLS embeddings.
 Brain diffuser [27] introduced a fine-grained representation of the image of the last hidden layer in CLIP.
@@
 Mind-Vis [5] conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values.
-Using a linear regression model, Takagi et al.
-[36] introduce additional ventral brain regions and align fMRI to the Stable Diffusion variable space.
-MindEye [33] uses a deep MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 [29] for the first time to narrow the disjointed vectors.
+MindEye [35] uses a large MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 [31] for the first time to narrow the disjointed vectors.
 At the same time, it creatively proposed a mixed contrast loss MixCo as a data augmentation method to make the large model converge.
+The recent trend in brain fMRI decoding research tends to use larger models to achieve better performance on downstream tasks [42, 16], ignoring the privacy and deployment issues discussed above.
 However, to the best of our knowledge, no previous research has focused on lightweight and efficient brain networks.
@@
 Fourier Transform plays a vital role in the area of digital signal processing.
-It has been introduced to deep learning for enhanced learning performance [11, 8, 43].
-GFNet [30] utilizes fast Fourier transform to convert images to the frequency domain and exchange global information between learnable filters.
+It has been introduced to deep learning for enhanced learning performance [11, 8, 45], and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis.
+In the field of computer vision, GFNet [32] utilizes Fast Fourier Transform(FFT) to convert images to the frequency domain and exchange global information between learnable filters.
 As a continuous global convolution independent of input resolution, Guibas et al.
-[12] design the adaptive Fourier neural operator frame token mixing.
+[12] design the Adaptive Fourier Neural Operator(AFNO) frame token mixing.
 Xu et al.
-[41] devise a learning-based frequency selection method to identify trivial frequency components and improve the accuracy of classifying images.
-On text classification, Lee-Thorp et al.
-[18] use the Fourier transform as a text token mixing mechanism.
-Furthermore, the Fourier transform is also applied to forecast time series  [3, 17, 16, 42].
+[43] devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy.
+As for natural language processing, Lee-Thorp et al.
+[20] use the Fourier Transform as a text token mixing mechanism on text classification.
+In research of time series forecast, Fourier Transform has deeper applications  [3, 18, 17, 44, 46, 19].
 To increase the accuracy of multivariate time-series forecasting, Cao et al.
-[3] propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the spectral domain.
+[3] propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain.
 Yang et al.
 [44] propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum.
-Wang et al.
-[38] prove that frequency domain MLP is more efficient than real domain MLP.
-fMRI is essentially a signal that can be processed in the frequency domain, but to the best of our knowledge, there has been no research on applying FFT to efficient brain representation.
-3 Preliminaries
-3.1 Motivation and Analysis
-FMRI signals are considered to have spatial redundancy, noise, and sample sparsity; even the nsdgeneral of NSD data has been preprocessed, which makes huge fully connected MLP of MindEye consume a lot of unnecessary parameter quantities.
-Furthermore, various data augmentation methods in MindEye are only used to prevent the overfitting of large models, which is also caused by MLP’s excessive focus on each voxel value.
-The fine-grained features come from the rich representation of the image by the last hidden layer of CLIP.
-Aligning multiple representations does not require excessive attention to each voxel’s values but rather to the different patterns of the entire brain region.
-The diffusion model that approximates disjointed cross-modal vectors is effective for retrieval.
-This paper does not discuss the lightweight of the diffusion model but rather discusses the efficient brain representation network backbone.
-We focus on designing a lightweight voxel encoder that can quickly capture global information of fMRI and efficiently align fMRI voxels into fine-gained image representations.
+Yi et al.
+[47] prove that frequency-domain MLP is more efficient than time-domain MLP.
+Processing fMRI frequency domain information is more helpful for understanding the mechanism of collaborative operation within the human brain.
+However, there is currently no research on processing fMRI signals in the frequency domain for representation learning.
+Figure 2: Overview of our Lite-Mind.
+Figure (a) is the architecture of the MLP Backbone of MindEye.
+Figure (b) is the architecture of the DFT backbone and Retrieval Pipeline.
+fMRI voxels are inputted into DFT Backbone to obtain voxel embeddings for downstream tasks.
 3 Lite-Mind
-Figure 2: Overview of our proposed Lite-Mind.
-Figure (a) is the architecture of the MLP Backbone of MindEye with 4 residual blocks(4096×4096).
-Figure (b) is our architecture of the DFT backbone and pipeline of retrieval.
-FMRI voxels are represented through DFT Backbone to obtain voxel embeddings.
-On the upper right is the overall retrieval pipeline, where LAION-5B retrieval and zero-shot classification are based on CLS token and diffusion projector.
-At the bottom right is the filter pyramid aggregation method, reducing the receptive field to align CLIP representation.
 3.1 Overview
-As illustrated in Figure 2, Lite-Mind comprises three components: 1) DFT Backbone maps flattened voxels to an intermediate space with a dimension of 257 × 768, corresponding to the last hidden layer of CLIP ViT/L-14.
-2) Retrieval pipeline contains various downstream retrieval-based tasks with the voxel embeddings from the DFT backbone.
-3) Contrastive Learning aims to bring the representation of two modalities closer together.
+As illustrated in Figure 2, Lite-Mind comprises two main components: 1) DFT Backbone consists of fMRI Spectrum Compression and fMRI Token Decoupling in Frequency-domain, mapping flattened voxels to an intermediate space.
+2) Retrieval pipeline contains Diffusion Projector, Contrastive Learning, and Downstream tasks with the voxel embeddings from the DFT backbone.
 3.2 DFT Backbone
-The backbone consists of patch partitioning followed by multilayer filter blocks, while the filter block adopts a pyramid aggregation strategy and finally outputs voxel embeddings.
-Given a fMRI-image sample pair [$x_{i}$,$y_{i}$] ($x_{i}$ represents the $i$-th fMRI sample, $y_{i}$ represents the image paired with it).
-Our optimization objectives are as follows:
+Sampling an fMRI-image pair ($x$,$y$) from dataset $D$, $x$ denotes the fMRI data and $y$ denotes the paired image.
+3.2.1 fMRI Spectrum Compression
+Patchify and Embedding The fMRI data $x$ is one-dimensional data composed of voxels.
+we firstly divide $x$ into $n$ non-overlapping patches $x=[x_{1},x_{2},...,x_{n}]$ using zero-padding, and adopt convolutions with positional encoding to obtain multiple tokens $t=[t_{1},t_{2},...,t_{n}]$.
+Filter Block. The spectrum of voxel tokens is obtained by 1D DFT as follows:
+$X[k]=F(t)=\sum_{i=1}^{n}t_{i}e^{-ki(2\pi/n)j}$ (1)
+Where $X\in\mathbb{C}^{n\times d}$ is a complex tensor, $X[k]$ is the spectrum of $t$ at the frequency $2\pi k/n$, $F(\cdot)$ denotes the 1D DFT along the voxel token dimension.
+$i$ denotes the $i$-th token, and $j$ denotes unit imaginary.
+After obtaining the spectrum of voxel tokens, we introduce learnable filters for filtering processing.
+Voxel spatial features are effectively consolidated within each frequency element, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain.
+We use $\textbf{K}=[\textbf{k}_{1},\textbf{k}_{2},...,\textbf{k}_{M}]$ to represent the filter library, where $M$ is the number of filters in the filter library.
+Spectrum $\hat{X}$ after filtering and compression is obtained as follows:
+$\hat{X}=\sum_{m=1}^{M}\frac{1}{n}{|X|}^{2}\odot\textbf{k}_{m}cos(\frac{(2m-1)\pi}{2M})$ (2)
+where $\odot$ is the element-wise multiplication, $|X|^{2}$ is the power spectrum of $X$.
+The $|X|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective.
+It also facilitates voxel information concentration and noise filtering.
+$cos((2m-1)\pi/2M)$ adopts the filterling form of Discrete Cosine Transform (DCT), which compacts better energy and can aggregate the more important information in fMRI voxels.
+Its combination of application with the channel filter library K allows for efficient frequency domain feature compression and noise denoising for fMRI.
+Finally, we employ Inverse Discrete Fourier Transform (IDFT), denoted by $F^{-1}(\cdot)$, to convert the spectrum back into the spatial domain:
+$\hat{t}\leftarrow F^{-1}(\hat{X})$ (3)
+3.2.2 fMRI Token Decoupling in Frequency-domain
+To align voxel tokens to image tokens, we decouple the filtered and compressed fMRI tokens again in frequency domain.
+Similarly, we follow Equation (2) to convert $\hat{t}$ to the frequency domain, then decouple tokens by a MLP-like model in frequency domain(FreMLP).
+FreMLP.
+Following Equation (2), the complex number output of $\hat{t}$ after DFT is $\hat{X}\in\mathbb{C}^{n\times d}$.
+Given a complex number weight matrix $\mathcal{W}\in\mathbb{C}^{n\times n^{\prime}}$, and a complex number bias $\mathcal{B}\in\mathbb{C}^{n^{\prime}}$, then the frequency-domain MLPs for token decoupling can be formulated as:
+$X^{\prime}={\sigma(\hat{X}^{T}\mathcal{W}+\mathcal{B})}^{T}$ (4)
+Where $X^{\prime}\in\mathbb{C}^{n^{\prime}\times d}$ is the final output, and $\sigma$ is the activation function.
+As both $\hat{X}$ and $\mathcal{W}$ are complex numbers, according to the rule of multiplication of complex numbers(see Appendix D.1 for details), we further extend the Equation (5) to:
+$\begin{split}X^{\prime}&=({\sigma(Re(\hat{X}^{T})\mathcal{W}_{r}-Im(\hat{X}^{T})\mathcal{W}_{i}+\mathcal{B}_{r})}\\ &+j\sigma(Re(\hat{X}^{T})\mathcal{W}_{i}+Im(\hat{X}^{T})\mathcal{W}_{r}+\mathcal{B}_{i}))^{T}\end{split}$ (5)
+Where $Re(\cdot)$ denotes the real part of $\hat{X}^{T}$, $Im(\cdot)$ denotes the imaginary part, $\mathcal{W}=\mathcal{W}_{r}+j\mathcal{W}_{i}$ and $\mathcal{B}=\mathcal{B}_{r}+j\mathcal{B}_{i}$.
+Due to the more significant features of filtered and compressed tokens in the frequency domain, we replace the final FC layer of MindEye’s MLP Backbone with FreMLP.
+For more theoretical proof and implementation of FreMLP, please refer to the Appendix D.2.
+Finally, we follow IDFT $t^{\prime}\leftarrow F^{-1}(X^{\prime})$ to get final tokens $t^{\prime}$, and use voxel embedding $f$ to represent them.
+3.3 Retrieval Pipeline
+3.3.1 Optimization Objective
+The optimization objective for the retrieval task is defined as follows:
 […]
-Patchfy.
-Given a fMRI-image pair $x_{i}=[v_{1},v_{2},...,v_{m}]$ ($i$ represents the $i$-th fMRI sample, $m$ is the length of fMRI voxels).
-we firstly divide $x_{i}$ into $n$ non-overlapping region patches $p_{i}=[p_{1},p_{2},...,p_{n}](m=nl)$ and adopt CNN to generate the multi-channel representation of voxel patches:
-$p^{c}_{i}=PE(CNN(x_{i}))$ (2)
-Where $PE(\cdot)$ is the region patch embedding for voxels, and $c$ is the number of patch channels.
-Filter Block. After we obtain the voxel patch embeddings $x_{i}=[p^{c}_{1},p^{c}_{2},...,p^{c}_{n}]$, then the spectrum of voxel patch embeddings is obtained by multi-channel 1D DFT as below:
-$X^{c}[k]=F_{pat}(x[i])=\sum_{i=0}^{n-1}x[i]e^{-j(2\pi/m)ki}$ (3)
-Where $X\in\mathbb{C}^{n\times d}$ is a complex tensor, $F_{pat}(\cdot)$ denotes the 1D DFT along the voxel patch dimension.
-Note that self-attention computes the spatial dependencies in a quadratic time complexity, while DFT can be efficiently implemented via a fast Fourier transform in logarithmic time complexity.
-After obtaining the spectrum of voxel patches, we use a voxel filter library for filtering processing.
-Voxel spatial features are effectively consolidated within each frequency element, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain.
-To expand voxel vectors to final hidden layers of CLIP with dimensions much larger, we introduce a filter library for each channel of voxel patches $X^{c}$, to obtain fine-grained alignment of images in the frequency domain.
-We use $K=[K_{c_{1}},K_{c_{2}},...,K_{c}]$ to represent the filter library, where c is the number of filters in the filter library:
-$\hat{X}=\sum_{i=1}^{c}\frac{1}{l}X^{i^{2}}\odot k_{i}cos(\frac{(2i-1)\pi}{2k})$ (4)
-where $\odot$ is the element-wise multiplication, $|X^{i}|^{2}$ is the power spectrum of $X^{i}$, $l$ is the length of $X^{i}$.
-The $|X^{i}|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective.
-It also facilitates voxel information concentration and noise filtering.
-$cos((2i-1)\pi/2k)$ compacts better energy and can aggregate the more important information in fMRI voxels.
-Its combination of application with the channel filter library $K$ allows for various fine-grained features alignment to images from different channel modes.
-Filter Pyramid Merge.
-To better align the voxel patches with the fine-grained representation of the image in the frequency domain, we adopt a filter pyramid merging strategy to expand the receptive field of the voxel patch channels.
-Gradually attenuated filter library converges to obtain final hidden layer $257\times 768$ for alignment, as shown in the lower right corner of Figure 2.
-4.3 Retrieval pipeline
+where $\omega$ is the weight of our DFT backbone, $SIM(\cdot)$ denotes the cosine similarity, and $CLIP(\cdot)$ denotes using CLIP model to extract image embedding.
+3.3.2 Diffusion Projector
+When searching on a large-scale dataset like LAION-5B, image-to-image performs better than fMRI-to-image and can find more similar images.
+Therefore, we used a diffusion projector to translate the disjointed voxel CLS embeddings of the DFT Backbone output, using a DALLE · 2[31] similar diffusion model to obtain the image CLS embeddings and retrieve it on LAION-5B: $\mathcal{V^{\prime}}=Diffusion(f)$.
+The Diffusion Projector consists of a transformer for image CLS embedding forecasting.
+3.3.3 Contrastive Learning
+Due to the inherent differences between different modalities, we use CLIP’s contrastive loss to train the DFT backbone.
+Sampling a batch $B$ from voxel-image pairs, the contrastive loss is defined below:
+$L_{contr}=-\frac{1}{|B|}\sum_{s=1}^{|B|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|B|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$ (7)
+where $f_{s}$ is the $s$-th voxel embeddings, $\mathcal{V}_{s}$ is its corresponding image embeddings, $\tau$ is a temperature factor.
+To perform image-to-image retrieval on LAION-5B, we use MSE loss to constrain the generation of approximate image embeddings:
+$L_{mse}=\frac{1}{|B|}\sum_{s=1}^{|B|}{||\mathcal{V}_{s}-\mathcal{V^{\prime}}_{s}||}^{2}_{2}$ (8)
+Accordingly, the training loss is defined below:
+$L=L_{contr}+\alpha L_{mse}$ (9)
+In general, when $\alpha=0$, Lite-Mind is trained to output voxel embeddings for retrieval on the NSD test set.
+However, for LAION-5B retrieval, $\alpha$ is non-zero to map the voxel CLS embeddings outputted by Lite-Mind to image CLS embeddings, enabling online retrieval.
+3.3.4 Downstream tasks
 The retrieval process on various downstream tasks is shown in the upper right of Figure 2.
-Test set retrieval. The cosine similarities between voxel embeddings and image embeddings are directly calculated.
-LAION-5B retrieval.
-Image embeddings are generated based on voxel CLS token by Diffusion Projector for online retrieval.
-Because our DFT backbone uses the same one-way contrast loss as CLIP, we consider that the DFT backbone outputs voxel embeddings similar to text embeddings, which we consider as voxel-level text.
-We found that when searching on the large-scale dataset LAION-5B, img2img performs better than text2img and can find more similar images.
-Therefore, we used a similar method to represent the disjointed voxels of the backbone output, using a DALLE · 2[29] similar diffusion model to obtain the image representation and retrieve it on LAION-5B.
+Test set retrieval. The cosine similarities between voxel embeddings from DFT Backbone and image embeddings from CLIP ViT are directly calculated.
+LAION-5B retrieval. Voxel CLS embeddings from the DFT backbone are translated to image CLS embeddings through the Diffusion Projector for online LAION-5B retrieval.
 Zero-shot classification.
-Image retrieval is performed on novel classes in the test set, and the similarities between the retrieved images and simple CLIP class text prompts are calculated for classification tasks.
-4.4 Contrastive Learning
-Due to the inherent differences between different modalities, we use CLIP’s contrastive loss to train the DFT backbone.
-To perform img2img retrieval on LAION-5B, we use MSE loss to constrain the generation of approximate image representations.
-Sampling a batch $\mathcal{B}$ from voxel-image pairs, the contrastive loss is defined below:
-$L_{contr}=-\frac{1}{|\mathcal{B}|}\sum_{s=1}^{|\mathcal{B}|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|\mathcal{B}|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$ (5)
-where $f_{s}$ is the $s$-th voxel representation, $\mathcal{V}_{s}$ is its corresponding image representation, $\tau$ is a temperature factor.
-Accordingly, the training objective is defined below:
-$L=\alpha L_{contr}+\beta L_{mse}$ (6)
-When training LAION-5B retrieval tasks, $\beta$ is a non-zero value; When training test set retrieval, $\beta=0$.
-Due to the lightweight performance of the model, we trained two separate models for fMRI-image retrieval and image-fMRI retrieval, corresponding to two types of one-way losses.
+Image retrieval is performed on novel classes on the test set, and the similarities between the retrieved images and simple CLIP class text prompts are calculated for classification tasks.
+Figure 3: Partial retrieval results of Lite-Mind on all 982 test images for Subject 1.
+With 12.5M DFT Backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates.
+The number below each image represents the similarity score.
+See more cases include failure retrieval in Appendix C.4.
 4 Experiments
@@
 For each test image, the fMRI responses from different trials were averaged.
-Figure 3: Partial search results of Lite-Mind on all 982 test images for Subject 1.
-Even without diffusion prior and heavy MLP backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates.
-The number below each image represents the similarity score.
+4.2 Implementation details
+To fairly compare different backbones, we did not use any data augmentation methods, such as voxel mixing loss MixCo and image slicing enhancement.
+Due to MindEye only disclosing the MLP Backbone performance of Subject 1, only the performance of Subject 1 is provided in Table 0(a).
+The experimental results of more subjects are shown in Appendix B.1.
+All of our experiments were conducted on a single Tesla V100 32GB GPU.
+More experimental details and hyperparameter settings can be found in Appendix A.
+5 Results
+5.1 fMRI/image retrieval
+Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings on the test set.
+If a paired image embedding is retrieved, the retrieval is considered correct.
+fMRI retrieval is the opposite process mentioned above.
+Note that there are many semantically and visually similar images on the NSD dataset test set, which are considered to be similar in CLIP space.
+Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image.
+For test retrieval, we adhered to the identical methodology as Lin et al.
+[21] and MindEye [35] to compute the retrieval metrics presented in Table 0(a).
+For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images.
+The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 982 voxel embeddings of the test set.
+We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples.
+The experimental results are shown in Table 0(a).
+We also conducted retrieval experiments on the remaining three subjects to demonstrate the universality of DFT Backbone.
+The detailed results and discussions are in Appendix B.2.
+As shown in Table 0(a), compared to MLP Backbone, our DFT Backbone improves retrieval accuracy by 5% and 14.9% for two retrieval ways, indicating that the fine-grained representation of the image comes from the rich representation of the last hidden layer of CLIP rather than MLP’s excessive attention to each voxel value.
+Note that our DFT Backbone is closest to MindEye’s MLP Backbone with only contrastive learning, while Prior in MLP Backbone+Prior includes MSE loss for image reconstruction.
 Method Model Parameters Retrieval
@@
 Lin et al. [21] deep models 2.34M 11.0% 49.0%
-Ozcelik… [25] 257 separate linear regression models 3B 21.1% 30.3%
-MindEye [33] MLP Backbone 940M 13.3% 63.1%
+Ozcelik… [27] 257 separate linear regression models 3B 29.9% 21.4%
+MindEye [35] MLP Backbone 940M 89.6% 82.2%
 MindEye [35] MLP Backbone+Projector 996M 88.8% 84.9%
-MindEye [33] MLP Backbone+Prior 3B 93.4% 90.1%
-Lite-Mind(ours) DFT Backbone 12.5M 94.3% 97.1%
+MindEye [35] MLP Backbone+Prior 1.2B 93.4% 90.1%
+Lite-Mind(ours) DFT Backbone 12.5M 94.6% 97.1%
 (a) Quantitative comparison of Lite-Mind retrieval performance against other models for Subject 1.
-Image retrieval refers to the percent of the time the correct image was retrieved out of 300 candidates, given the associated brain sample (chance=0.3%); vice-versa for brain retrieval.
-Lite-Mind only uses a parameter quantity of 12.5M, achieving extremely high retrieval performance without using any model to close vectors of different modalities(see Appendix B for remaining subject models).
+Image retrieval refers to the hit rate of correct retrieval from 300 candidates, given the associated brain sample (chance=0.3%); vice-versa for brain retrieval.
+Lite-Mind only uses a parameter quantity of 12.5M, achieving extremely high retrieval performance without using any model to close vectors of different modalities(see Appendix B.1 for remaining subject models).
 Method Low-Level High-Level
@@
 All results are averaged across four subjects, and Lite-Mind achieved higher performance with a lighter backbone, even exceeding the accuracy of many reconstruction methods.
-5.2 Implementation details
-To fairly compare different backbones, we did not use any data augmentation methods, such as voxel mixing loss MixCo and image slicing enhancement.
-Due to MindEye only disclosing the MLP Backbone performance of Subject 1, only the performance of Subject 1 is provided in Table 0(a).
-The experimental results of more subjects are shown in Appendix B.
-All of our experiments were conducted on a single Tesla V100 32GB GPU.
-More experimental details and hyperparameter settings can be found in Appendix A.
-6 Results
-6.1 fMRI/image retrieval
-Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings in the training set.
-If a paired image embedding is retrieved, the retrieval is considered correct.
-fMRI retrieval is the opposite process mentioned above.
-Note that there are many semantically and visually similar images in the NSD dataset test set, which are considered to be similar in CLIP space.
-Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image.
-For test retrieval, we adhered to the identical methodology as Lin et al.
-[19] and MindEye [33] to compute the retrieval metrics presented in Table 0(a).
-For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images.
-The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 982 voxel embeddings of the test set.
-We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples.
-The experimental results are shown in Table 0(a).
-We also conducted retrieval experiments on the remaining three subjects to demonstrate the universality of DFT Backbone.
-The detailed results and discussions are in Appendix B.
-As shown in Table 0(a), compared to MLP Backbone+Projector our DFT Backbone improves retrieval accuracy by 5.5% and 12.2% for two retrieval ways, indicating that the fine-grained representation of the image comes from the rich representation of the last hidden layer of CLIP rather than MLP’s excessive attention to each voxel value.
-Furthermore, our DFT Backbone retrieval performance even surpasses MindEye’s MLP Backbone+Prior without any image enhancement methods, proving that frequency domain calculations better capture global features.
-Figure 4: Partial search results of Lite-Mind on LAION-5B retrieval for Subject 1.
-The left column marked by a red box in every two columns represents the original image seen by the subject, and the right column represents the image retrieved by LAION-5B.
+Table 1: Main retrieval results on the NSD and LAION-5B dataset.
+5.2 LAION-5B retrieval
+Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B [34].
+The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client [2].
+For LAION-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS embeddings of CLIP.
+For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS embeddings, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP).
+Unlike MindEye, in the LAION-5B retrieval process, we used the Diffusion Projector to conduct image-to-image retrieval.
+To compare the performance of large-scale dataset retrieval, we use metrics of image reconstruction for evaluation with previous image reconstruction methods.
+The experimental results are shown in Table 0(b).
+The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric [40]; EfficientNetB1(”Eff”) [39] and SwAV-ResNet50(”SwAV”) [4] refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
+Experiments show that our DFT Backbone performs better than MindEye in high-level metrics, as shown in Table 0(b), although our backbone only has 8M parameters(MindEye-CLS backbone still has 125M.
+Note that both MindEye and Lite-Mind perform LAION-5B retrieval based on CLS alignment, resulting in a smaller number of parameters compared to full models).
+The results indicate that an efficient DFT backbone can also assist downstream retrieval tasks.
+Partial retrieval visualization results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval and visualization results of each Subject are shown in Appendix B.2.
+5.3 GOD zero-shot classification
+Since the collective fMRI data are much fewer and shorter than those of the NSD dataset usually, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify the Lite-Mind’s generalization ability.
+The voxel lengths collected by Subjects on the GOD dataset range from 4133 to 4643, approximately 30% of NSD, and there are only 1200 fMRI-image pairs on the training set (4.8% of NSD).
+We conduct zero-shot classification by using fMRI-to-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification).
+We train our Lite-Mind-15.5M to align voxels to images on the GOD dataset.
+We also replicated the MLP backbone and projector of MindEye, changing the residual block of $4096\times 4096$ to $1024\times 1024$ to accommodate the reduction of voxel length(We empirically found that the original $4096\times 4096$ performed even worse), and MindEye still has 210M parameters.
+As shown in Table 2, the experimental results indicate that our proposed Lite-Mind still guarantees robustness and establishes a new state-of-the-art for zero-shot classification on the GOD dataset without complex text prompt templates or auxiliary text representation training.
+On the contrary, due to the sudden decrease in dataset size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data size for voxel value-wise mappings with large MLP Backbone.
+More retrieval and classification results are in Appendix B.3.
+Figure 4: Partial retrieval results of Lite-Mind from LAION-5B for Subject 1.
+The left column marked by a red box in every two columns represents the original image seen by the Subject, and the right column represents the image retrieved from LAION-5B.
 Methods Modality Prompt Subject1 Subject2 Subject3 Subject4 Subject5
@@
 Lite-Mind-15.5M(ours) V Text 26.00 74.00 28.00 70.00 30.00 80.00 34.00 76.00 24.00 66.00
-Table 1: Zero-shot visual stimulus classification on GOD dataset.
+Table 2: Zero-shot visual stimulus classification on the GOD dataset.
 The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%).
 The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from [7].
-MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B for retrieval performance on the GOD dataset).
-6.2 LAION-5B retrieval
-Because Lite-Mind can also align image representation space, we can expand the retrieval task to larger datasets, such as LAION-5B [32].
-For LIAON-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS token of CLIP.
-The final layer CLIP ViT-L/14 embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client [2].
-For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS token, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP).
-Unlike MindEye, in the LAION-5B retrieval process, we used the diffusion model of DALLE·2 prior to conducting img2img retrieval.
-To compare the ability of LAION-5B retrieval to replace image reconstruction, we compared the performance indicators of reconstruction with previous image reconstruction methods.
-The experimental results are shown in Table 0(b).
-The specific evaluation indicators in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstructions; SSIM is structural similarity index metric [39]; EfficientNetB1(”Eff”) [37] and SwAV-ResNet50(”SwAV”) [4] refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
-Experiments show that our diffusion projector performs better than the projector of MindEye in high-level metrics, as shown in Table 0(b), although our backbone only has a parameter size of 8M (MindEye-CLS backbone still has 125M).
-The results indicate that efficient DFT backbones can also assist downstream retrieval tasks.
-Partial search results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval performance and results of each subject are shown in Appendix B.
-6.3 GOD zero-shot classification
-Since the collective prime values and data volume of fMRI data are much smaller than those of the NSD dataset in most cases, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify the Lite-Mind’s generalization ability.
-The voxel values collected by participants in the GOD dataset range from 4133 to 4643, much smaller than that of the NSD dataset, and the training set has only 1200 fMRI-image pairs(4.8% of NSD training set).
-We conduct zero-shot classification by using fMRI-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification).
-We train our Lite-Mind-15.5M to align voxels to images on the GOD dataset.
-We also replicated the MLP backbone and projector methods of MindEye, changing the residual block of $4096\times 4096$ to $1024\times 1024$ to accommodate the reduction in voxel vector length, and MindEye still has a parameter quantity of 210M.
-As shown in Table 1, the experimental results indicate that our proposed Lite-Mind still guarantees versatility and establishes a new state-of-the-art for zero-shot classification on the GOD dataset without complex text prompt templates or auxiliary text representation training.
-On the contrary, due to the sudden decrease in data size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data volume and data augmentation for voxel value-wise mappings in heavy MLP Backbone.
-More retrieval and classification results are provided in Appendix B.
+MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B.3 for retrieval performance on the GOD dataset).
 5.4 Ablations and visualization
+We assess the effectiveness of different modules within Lite-Mind to investigate the lightweight and efficient performance in this session. All experiments below are for Subject 1 on the NSD dataset.
 Architectural Analysis.
-We assess the effectiveness of different modules within Lite-Mind to investigate the lightweight and efficient performance in this session.
-All ablation experiments are for Subject 1 on the NSD dataset.
-We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of DFT on retrieval efficiency(Table 2).
-We even completely removed the DFT layer and aligned the patches directly to the last hidden layer of CLIP ViT-L/14 with a tiny MLP, still achieving a forward retrieval accuracy of 65.8%, indicating that patches have more concentrated information than voxels.
-As the number of DFT layers gradually deepens, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual.
-We interestingly found that the accuracy of brain retrieval tends to converge faster than the image and only 6 layers of DFT are needed to approach the highest value.
-We have also conducted more ablation experiments on data volume to demonstrate the robustness of Lite-Mind.
-Refer to Appendix B for more ablations.
+We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of fMRI Spectrum Compression on retrieval efficiency(Table 3).
+As the Filter Blocks layers gradually deepen, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual.
+We interestingly found that the accuracy of brain retrieval tends to converge faster than the image retrieval and only 6 layers of Filter Blocks are needed to approach the highest value.
+We have also conducted more ablation experiments on hyper-parameters to demonstrate the robustness of Lite-Mind.
+Refer to Appendix B.4 for it.
+Depth FLOPs(G) Parameters Image Retrieval Brain Retrieval
+MindEye 5.66 996M 88.8% 84.9%
+1 0.05 0.7M 83.5% 94.0%
+6 0.22 3.4M 90.6% 97.2%
+12 0.43 6.6M 91.0% 97.2%
+21 0.80 12.5M 94.6% 97.4%
+Table 3: Ablation experiments on fMRI Spectrum Compression revealed a positive correlation between the depth of the Filter Blocks and enhanced forward retrieval accuracy.
+Module Analysis.
+We conducted ablation experiments on two main modules of DFT Backbone in Table 4.
+We completely removed the Filter blocks and aligned tokens directly to the image embeddings with fMRI token decoupling, achieving a forward retrieval accuracy of 65.8%, indicating that the importance of frequency domain filtering for denoising and compression.
+Then we replaced FreMLP with the FC layer in the real domain, where tokens are flatten into a tensor for alignment, similar to the last layer of the MLP backbone of MindEye.
+The model performance rapidly declined, indicating that tokens processed by DFT Backbone have better mapping performance with FreMLP in the frequency domain.
+Module Image Retrieval Brain Retrieval
+DFT Backbone 94.6% 97.4%
+w/o Filter Blocks 65.8% 74.7%
+w/o FreMLP 43.4% 47.8%
+Table 4: Ablation experiments on modules of DFT Backbone.
+Embedding Dimension.We conducted additional experiments to verify the impact of embedding dimensions on retrieval accuracy, as shown in Table 5.
+Among them, two different CLIP models (i.e., CLIP ViT/B-32 and CLIP ViT/L-14) were used to extract image embeddings, with a total of four embedding dimensions, by CLS embeddings and the last hidden layer, respectively.
+From the results, it can be seen that the retrieval accuracy of Lite-Mind is higher when the embedding dimensions of the image are longer and the representation of the image is richer.
+The results verify that the fine-grained alignment we mentioned above comes from the rich representation of the image, rather than the fMRI voxel value fully connected heavy MLP Backbone.
+Embeddings Dimension Parameters Retrieval
+[…]
+ViT-B/32 CLS 512 6.7M 57.5% 61.8%
+ViT-L/14 CLS 768 6.7M 60.3% 64.2%
+ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
+ViT-L/14 Hidden 257×768 12.5M 94.6% 97.4%
+Table 5: Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
+Cerebral cortex.
+In order to explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s [38] method to extract the stream ROI on the NSD dataset, covering the visual cortex of nsdgenal, and trained the individual retrieval DFT Backbones for visual cortices (early, lateral, parietal, ventral).
+As shown in Table 6, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy.
+Although with voxel length of 5917, Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research.
+They demonstrated that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving fully connected MLP Backbones are unnecessary.
+Interestingly, for other cortical regions, both lateral and ventral have some impacts on retrieval accuracy, while parietal shows little.
+Region Voxel Length Parameters Retrieval
+[…]
+early 5917 12.4M 85.0% 93.4%
+lateral 7799 12.4M 26.9% 29.0%
+parietal 3548 12.3M 9.6% 11.2%
+ventral 7604 12.3M 18.1% 22.4%
+Table 6: Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
 Visualization.
-We visualize the diffusion model representation of LAION-5B using T-SNE in Figure 5.
-The graph shows that the diffusion model successfully transformed the voxel representation learned from contrastive learning into an image representation, and the two representations were well fused together for img2img retrieval on the LAION-5B dataset.
-We have also conducted more visualization experiments in Appendix C to demonstrate the efficiency of DFT Backbone.
-DFT depth Parameters Image Retrieval Brain Retrieval
-Pyramid(18) 12.5M 94.3% 97.4%
-18 9.9M 91.8% 97.3%
-12 6.6M 91.0% 97.2%
-6 3.4M 90.6% 97.2%
-1 0.7M 83.5% 94.0%
-0 0.2M 65.8% 74.7%
-Table 2: Ablation experiments on Discrete Fourier Transform (DFT) layers revealed a positive correlation between the depth of the DFT layer and enhanced forward retrieval accuracy.
-Simultaneously, augmenting the filter pyramid of the receptive field demonstrated an additional improvement in retrieval accuracy, provided there is a judicious increase in the parameter count.
-Figure 5: T-SNE of representations of 982 test fMRI voxels by diffusion projector on LAION-5B retrieval for Subject 5.
-The diffusion model plays a role in bringing vectors closer, enabling img2img retrieval on the LAION-5B dataset.
+We visualize the CLS embeddings for LAION-5B using T-SNE in Figure 5.
+The figure shows that the diffusion projector successfully transformed the voxel CLS embeddings learned from contrastive learning into the image CLS embeddings, and the two kinds of embeddingss are well fused together for image-to-image retrieval on the LAION-5B dataset.
+We have also visualized the alignment process of voxel embeddings by contrastive learning on the NSD dataset in Appendix C.3, and learned filters of Filter Blocks in fMRI Spectrum Compression in Appendix C.5.
+Figure 5: T-SNE of embeddings of 982 test fMRI by diffusion projector on LAION-5B retrieval for Subject 1.
+In the figure, image-clip represents image CLIP CLS embeddings, while voxel-fft and voxel-diffusion represents voxel CLS embeddings by DFT Backbone or Diffusion Projector.
+The diffusion projector plays a role in bringing vectors closer, enabling image-to-image retrieval on the LAION-5B dataset.
+6 Limitations and Future Work
+Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models.
+Although our exploratory experiments have shown that a common model can achieve comparable retrieval accuracy on four Subjects, it’s still necessary to ensure sufficient training data.
+Further design beyond the model level is still needed for cross-Subject models, to achieve Subject-specific few-shot or even zero-shot.
+In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence.
+This provides a promising avenue for the interpretability of deep learning in neuroscience.
 7 Conclusion
-In this work, we propose Lite-Mind, an extremely lightweight brain representation extraction network based on Fourier transform for cross-modal fMRI-image retrieval.
+In this work, we propose Lite-Mind, an extremely lightweight brain representation learning paradigm based on Fourier Transform for cross-modal fMRI-to-image retrieval.
 Our DFT Backbone is an efficient means of obtaining fine-grained representation alignment between fMRI signals and visual stimuli.
 With high retrieval accuracy, Lite-Mind can transform many downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification.
-Meanwhile, Lite-Mind’s dependence on data volume and voxel values is less than that of larger models, demonstrating excellent universality across different fMRI datasets.
-For future work, we should explore the relationship between different sizes of patches and brain area structures and explore the impact of cortices other than the visual cortex, such as ventral, on retrieval accuracy.
-In addition, versatile cross-subject models should be studied to efficiently project fMRI representations of different subjects into the same joint space or transfer from subject models with larger sample sizes to models with sparse sample sizes.
-At the same time, how to use an efficient model to collaboratively train multiple fMRI datasets (across subjects) is also a promising direction that can be explored in the future.
+Meanwhile, Lite-Mind’s dependence on dataset size and voxel length is less than larger models, demonstrating excellent robustness across different fMRI datasets.
 Supplementary Material
@@
 Hyper-parameters.
-On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and $\beta=0$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval.
-Owning to the lightweight nature of Lite-Mind, the batch size is set to 1500, and the learning rate is 1.16e-3.
-For Subject 1, 2, 5, 7, Pyramid is set to [512, 256, 128, 257] respectively, and the corresponding number of layers and the patch size are set to [2, 10, 2, 4] and [480, 440, 460, 450], respectively.
+On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval.
+Owning to the lightweight nature of Lite-Mind, the batch size is set to 500, the learning rate is 1e-3, patch size is set to 480, and Filter library size is 4.
+Filter block layers are the same for all subjects.
 About LAION-5B retrieval, for DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, while the weight decay is 6.02e-2 for diffusion projector.
-CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, and $\beta=0.5$, the batch size is set to 80 and the learning rate is 1.16e-3 for DFT Backbone while 1.1e-4 for diffusion projector.
+CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, the batch size is set to 80 and the learning rate is 1e-3 for DFT Backbone while 1.1e-4 for diffusion projector.
+Note that hyper-parameters in Diffusion Projector are the value of the open-source DALLE·2 from https://github.com/lucidrains/DALLE2-pytorch.
 On the GOD dataset, the hyperparameters are the same as those on the NSD dataset, except for the batch size and patch size which are set to 1200 and 8 for all 5 subjects on the GOD dataset, respectively.
@@
 To demonstrate the applicability of our method, we also conducted experiments on three other Subjects, i.e., Subject 2, Subject 5, and Subject 7, on the NSD dataset, and the experimental results are shown in Table 7.
-In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths, but instead used smaller patches.
-It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90, and the retrieval accuracies of Subjects 5 and 7 are also above 75, The results prove that DFT Backbone can efficiently work on different subjects.
+In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths.
+It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90%, and the retrieval accuracies of Subjects 5 and 7 are also above 80%.
+The results prove that DFT Backbone can efficiently work on different subjects.
 Note that MindEye only includes the results of the overall model in Subjects 2, 5, and 7 in its Appendix, and it does not evaluate the effect of individual MLP Backbone.
@@
 […]
-Lite-Mind(Subj 1) 15724 12.51M 94.3% 97.4%
-Lite-Mind(Subj 2) 14278 12.49M 93.4% 98.2%
+Lite-Mind(Subj 1) 15724 12.51M 94.6% 97.4%
+Lite-Mind(Subj 2) 14278 12.49M 94.1% 98.2%
 Lite-Mind(Subj 5) 13039 12.47M 80.5% 86.3%
-Lite-Mind(Subj 7) 12682 12.47M 75.0% 82.3%
+Lite-Mind(Subj 7) 12682 12.47M 81.7% 82.3%
 Table 7: Addtional retrieval performance for individual subjects on 982 test images of the NSD dataset.
@@
 Meanwhile, as shown in Figure 7, the retrieval performance of LAION-5B completely depends on the retrieval accuracy of the CLS model in the test set, such that images retrieved incorrectly in the test set may also have retrieval bias on LAION-5B, for example treating a teddy bear as an image of a cat or dog as shown in Figure 7.
-However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS token models.
-In the future, it would be beneficial to explore models that improve the alignment of CLS tokens or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
+However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS embedding models.
+In the future, it would be beneficial to explore models that improve the alignment of CLS embeddings or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
 Figure 7: Addtional retrieval results corresponding to Figure 4.
@@
 Table 9: Retrieval performance on the GOD dataset.
-B.4 Additional Ablations
-In this section, we delved into the dependence of Lite-Mind on data volume and feature length on the NSD dataset and explored the specific impact of different cerebral cortexes on retrieval accuracy, not just nsdgenal.
-All the experimental results on the NSD dataset are derived from Subject 1, with a retrieving pool size of 300.
-B.4.1 Retrieval with varying dataset size
-In order to better investigate the versatile representation of DFT Backbone, we conducted data volume experiments on the NSD dataset, as shown in Table 6.
-Among them, All data corresponds to all trials of every single Subject on the NSD dataset, and 24890 fMRI-image pairs are used for training; Half data refers to randomly removing half of the training data; 2-session refers to using only 1500 randomly selected fMRI-image pairs for training, which corresponds to the two scan sessions of the fMRI experiment.
-In the table, we report the results of MLP Backbone and diffusion models for MindEye.
-The experimental results indicate that even in the same NSD dataset, the dependence of Lite-Mind on the number of training samples is smaller than that of MindEye.
-This also proves that the lightweight model has a wider applicability, i.e., versatility.
-Method Parameters Retrieval
+B.4 Hyper-parameters Experiment
+In this section, we explore the influence of some hyperparameters on the model, including patch size and number of filters $M$, to verify the parameter sensitivity of the model.
+All the experimental results on the NSD dataset are from Subject 1, with a retrieving pool size of 300.
+Patch size.
+We conducted experiments by varying the patch size as presented in Table 10.
+The results exhibit stable retrieval accuracy above 90%, which indicates that our DFT Backbone is not sensitive to part size.
+Since we found patch size has less effect on retrieval accuracy, we chose a relatively large patch size (e.g., 480) to ensure fewer parameters and faster convergence.
+Patch size Parameters Retrieval
 […]
-All Data(MindEye) 3B 97.2% 94.7%
-Half Data(MindEye) 3B 77.5% 60.8%
-2-Sessions(MindEye) 3B 17.9% 12.0%
-All Data(Lite-Mind) 12.5M 94.3% 97.4%
-Half Data(Lite-Mind) 12.5M 83.4% 91.9%
-2-Sessions(Lite-Mind) 12.5M 11.9% 14.1%
-Table 6: Retrieval performance with varying dataset size for Subject 1 on the NSD dataset.
+50 14.0M 91.6% 96.4%
+200 12.7M 92.9% 96.9%
+480 12.5M 94.6% 97.1%
+600 12.3M 94.2% 97.0%
+900 11.9M 90.1% 95.5%
+Table 10: Retrieval performance for different patch size on the NSD dataset.
+Filter library size.
+We conducted experiments by varying the Filter library size as presented in Table 11.
+Specifically, there is a significant performance improvement from $M=1$ to $M=4$, while a slight fluctuation is observed for $M=4/8$.
+By setting $M=4$, the model has the ability to acquire diverse and distinct feature patterns from various dimensions of the frequency response while still maintaining an appropriate computational cost.
+Therefore, we determine that $M=4$ is the optimal choice on the NSD dataset.
+Filters Parameters Retrieval
+[…]
+1 9.0M 89.8% 96.4%
+2 10.2M 93.5% 97.4%
+4 12.5M 94.6% 97.1%
+8 17.2M 94.4% 96.7%
+Table 11: Retrieval performance for different Filter library size on the NSD dataset.
 Figure 8: Lite-Mind’s training loss curve and testing accuracy curve for Suject 1, 2, 5, 7 on the NSD dataset. The testing accuracy is calculated based on a retrieval pool of size 982.
-B.4.2 Retrieval with varying representation sizes
-We conducted additional experiments to verify the impact of representation dimensions on retrieval accuracy, as shown in Table 7.
-Among them, two different CLIP models (i.e., CLIP ViT/B-32 and CLIP ViT/L-14) were used to extract image representation, with a total of four representation dimensions, by CLS token and the last hidden layer, respectively.
-From the results, it can be seen that the retrieval accuracy of Lite-Mind is higher when the representation dimension of the image is longer and the representation of the image is richer.
-The results verify that the fine-grained alignment we mentioned in the main text comes from the rich representation of the image, rather than the fMRI voxel value fully connected heavy MLP Backbone.
-Method Representation Parameters Retrieval
-[…]
-ViT-B/32 CLS 512 6.7M 57.5% 61.8%
-ViT-L/14 CLS 768 6.7M 60.3% 64.2%
-ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
-ViT-L/14 Hidden 257×768 12.5M 94.3% 97.4%
-Table 7: Retrieval performance with different CLIP representation for Subject 1 on the NSD dataset.
-B.4.3 Cerebral cortex
-In order to explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s method to extract the stream region of the NSD dataset, covering the visual cortex of nsdgenal, and separately trained the retrieval model for each cortex (early, lateral, parietal, ventral).
-As shown in Table 8, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy, although with only 5917 voxels Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research.
-They believe that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving that fully connected backbones are unnecessary.
-Interestingly, for other cortical regions, both lateral and ventral have a good impact on retrieval accuracy, while parietal shows little impact.
-The findings indicate the necessity of further research on whether the accuracy improvement is related to the longer voxel length of these two regions.
-Region Voxel Length Parameters Retrieval
-[…]
-early 5917 12.4M 85.0% 93.4%
-lateral 7799 12.4M 26.9% 29.0%
-parietal 3548 12.3M 9.6% 11.2%
-ventral 7604 12.3M 18.1% 22.4%
-Table 8: Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
 Figure 9: Lite-Mind’s retrieval heatmap on the NSD dataset for Subject 1.
@@
 It shows that Lite-Mind has effectively retrieved corresponding images, even if there are many similar images in the test set, which verifies the fine-grained ability of Lite-Mind.
-C.3 Visualization in Frequency Domain
-We visualized the weights of different patches in different layers of filters, as shown in Figure 10 and 11.
-Visualization is divided into the real part (left) and imaginary part (right) of filter weights.
-It can be observed that different channels have varying degrees of attention to different patches, and the frequency domain better captures this characteristic.
-Interestingly, the weight of the imaginary part for the first and last patches is almost always 0, indicating that the noise is distributed in these two patches.
-(a) The real part and imaginary of patch weights for CLS token channel.
-(b) The real part and imaginary of patch weights for 140th channel.
-Figure 10: Weights visualization for DFT Backbone of $257\times 768$ representation length.
-(a) The real part and imaginary of patch weights for the 1st channel.
-(b) The real part and imaginary of patch weights for the 60th channel.
-Figure 11: Weights visualization for DFT Backbone of 768 representation length.
-C.4 Information Alignment
+C.3 Information Alignment
 We visualized the T-SNE plot between the voxel embeddings output by DFT Backbone and the image embeddings of frozen CLIP as the accuracy of the test set improved, as shown in Figure 13.
 We can observe that as the training progresses, the retrieval accuracy of the test set improves, and the shape of voxel embeddings tends to be closer to image embeddings, indicating the success of contrastive learning.
-Appendix D Efficiency Discussion
-D.1 Theoretical Analysis.
-Theorem.
-Suppose that H is the representation of raw fMRI voxel patches and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of a voxel patch series in the spatial domain is equal to the energy of its representation in the frequency domain.
+C.4 More Retrieval Cases
+In this section, we visualize retrieval failure cases in Figure 10, although only dozens of images are not in the Top 1.
+From the left half of the Figure, it can be seen that though not in Top 1, Lite-Mind still retrieved ground-truth in Top 2, and images in Top 4 are similar(either semantically similar or structurally similar, eg.
+animals in the wild or an airplane on the runway).
+On the contrary, a smaller proportion of ground-truth did not appear within the Top 4, as shown in the right half of the Figure.
+From the perspective of the images themselves, most of the scenes are too complex, even abstract (as shown in the second image), which may be the reason for the retrieval failure.
+Figure 10: Partial failure retrieval results of Lite-Mind on all 982 test images for Subject 1. The number below each image represents the similarity score.
+(a) The real part and imaginary of a filter weights.
+(b) The real part and imaginary of another filter weights from the same Filter Library.
+Figure 11: Weights visualization for DFT Backbone of $257\times 768$ embedding length.
+(a) The real part and imaginary of a filter weights.
+(b) The real part and imaginary of another filter weights from the same Filter Library.
+Figure 12: Weights visualization for DFT Backbone of 768 CLS embedding length.
+C.5 Visualization in Frequency Domain
+We visualized the weights of Filter library in Filter Blocks, as shown in Figure 11 and 12.
+Visualization is divided into the real part (left) and imaginary part (right) of filter weights.
+It can be observed that different filters have varying degrees of attention to different tokens, and the frequency domain better captures this characteristic.
+Interestingly, the weight of the imaginary part for the first and last tokens is almost always 0, indicating that the noise is distributed in these two tokens.
+Appendix D Theoretical Analysis.
+D.1 Complex Multiplication.
+For two complex number values $\mathcal{Z}_{1}=(a+jb)$ and $\mathcal{Z}_{2}=(c+jd)$, where $a$ and $c$ is the real part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively, $b$ and $d$ is the imaginary part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively.
+Then the multiplication of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ is calculated by:
+$\mathcal{Z}_{1}\mathcal{Z}_{2}$ $=(a+jb)(c+jd)$ (10)
+$=ac+j^{2}bd+jad+jbc$
+$=(ac-bd)+j(ad+bc)$
+D.2 Theorem Proof.
+Theorem 1.
+Suppose that H is the representation of raw fMRI voxel tokens and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of voxel tokens in the spatial domain is equal to the energy of its representation in the frequency domain.
 Formally, we can express this with the above notations:
 $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$ (11)
-Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the spatial/channel dimension, $f$ is the frequency dimension.
+Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the token dimension, $f$ is the frequency dimension.
 Proof.
-Given the representation of raw voxel patch series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in either the $C$ dimension (channel dimension) or the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
+Given the representation of raw voxel token series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
 $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{\textbf{H}(v)}{\textbf{H}^{*}(v)}\,\mathrm{d}v$ (12)
-where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
-$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v$ $=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v$ (9)
-=*$=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f$
-=*$=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f$
-$=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$
+where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}\textbf{}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
+$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v&=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v\\ &=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f\\ &=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f\\ &=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f\end{aligned}$
 Proved. ∎
-Therefore, the energy of a voxel patch series in the spatial domain is equal to the energy of its representation in the frequency domain.
-Complexity Analysis.
-For a fMRI voxel with a length of $m$, we divide it into $n$ $l$-length patches ($m=nl$) with $C$ channel numbers.
-Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP and DFT respectively, the middle layer dimension of MLP is $D_{1}$, and the alignment representation dimension is $N\times D_{2}$, where $N$ is the number of the last hidden layer of CLIP.
-The time complexity of MLP Backbone is $O(mD_{1}+L_{1}D_{1}^{2}+ND_{1}D_{2})$.
-For DFT Backobone, the time complexity of patchfy is $O(mC)$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nC\operatorname{log}n+nC)$.
-The time complexity of each layer’s tiny MLP is $O(8nC^{2})$.
+Therefore, the energy of a voxel token series in the spatial domain is equal to the energy of its representation in the frequency domain.
+Theorem 2.
+Given the voxel token series input H and its corresponding frequency domain conversion H, the operations of frequency-domain MLP on H can be represented as global convolutions on H in the spatial domain.
+This can be given by:
+$\mathcal{H}\mathcal{W}+\mathcal{B}=\mathcal{F}(\textbf{H}*W+B)$ (14)
+Where $\mathcal{F}$ is DFT, *** is a circular convolution, $\mathcal{W}$ and $\mathcal{B}$ are the complex number weight and bias, and $W$ and $B$ are the weight and bias in the spatial domain.
+Proof.
+Suppose that we conduct operations in the N (i.e., token dimension), then
+$\mathcal{F}(\textbf{H}(v)*W(v))=\int_{-\infty}^{\infty}{(\textbf{H}(v)*W(v))e^{-j2\pi fv}}\,\mathrm{d}v$ (15)
+According to convolution theorem, $\textbf{H}(v)*W(v)=\int_{-\infty}^{\infty}(\textbf{H}(\tau)W(v-\tau))\,\mathrm{d}\tau$, then
+$\mathcal{F}(\textbf{H}(v)*W(v))&=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{(\textbf{H}(\tau)W(v-\tau))e^{-j2\pi fv}}\,\mathrm{d}\tau}\,\mathrm{d}v\\ &=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(v-\tau)e^{-j2\pi fv}}\,\mathrm{d}v}\textbf{H}(\tau)\,\mathrm{d}\tau\end{aligned}$
+Let $x=v-\tau$,then
+$\mathcal{F}(\textbf{H}(v)*W(v))&=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi f(x+\tau)}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi fx}e^{-j2\pi f\tau}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{\infty}{(\textbf{H}(\tau)*)e^{-j2\pi f\tau}}\,\mathrm{d}\tau\int_{-\infty}^{\infty}{(\textbf{W}(x)*)e^{-j2\pi fx}}\,\mathrm{d}x\\ &=\mathcal{H}(f)\mathcal{W}(f)\end{aligned}$
+Accordingly, $\textbf{H}(v)*W(v)$ in the spatial domain is equal to $\mathcal{H}(f)\mathcal{W}(f)$ in the frequency domain.
+Therefore, the operations of FreMLP ($\mathcal{H}\mathcal{W}+\mathcal{B}$) in the token dimension (i.e., v = N ) are equal to the operations ($\mathcal{F}(\textbf{H}*W+B)$) in the spatial domain.
+This implies that frequency-domain MLPs can be viewed as global convolutions in the spatial domain.
+Proved.
+∎
+D.3 Complexity Analysis
+For a fMRI voxel with a length of $l$, we divide it into $n$ patches.
+Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP Backbone and DFT Backbone respectively, the middle layer dimension of MLP Backbone is $D$, and the alignment embedding dimension is $n^{\prime}\times D^{\prime}$, where $n^{\prime}$ is the the number of tokens of CLIP.
+The time complexity of MLP Backbone is $O(lD+L_{1}D^{2}+n^{\prime}DD^{\prime})$.
+For DFT Backobone, the time complexity of patchify and tokenization is $O(lD^{\prime})$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nD^{\prime}\operatorname{log}n+nD^{\prime})$.
+The time complexity of FreMLP is $O(2nD^{\prime}\operatorname{log}n+2nn^{\prime}D^{\prime}+2n^{\prime}D^{\prime})$.
 Thus the time complexity of the entire DFT Backbone is:
-$O(mC+L_{2}(2nC\operatorname{log}n+nC+8nC^{2})+nD_{2})$ (10)
-$=O(mC+nC\operatorname{log}nL_{2}+nC^{2}L_{2}+nD_{2})$
-Quantitative analysis algorithm complexity for DFT Backbone is shown in Table 9.
-Model FLOPs(G) Parameters Image Retrieval Brain Retrieval
-MindEye 5.66 996M 88.8% 84.9%
-Pyramid-18 0.80 12.5M 94.3% 97.4%
-DFT-18 0.63 9.9M 91.8% 97.3%
-DFT-12 0.43 6.6M 91.0% 97.2%
-DFT-6 0.22 3.4M 90.6% 97.2%
-DFT-1 0.05 0.7M 83.5% 94.0%
-DFT-0 0.02 0.2M 65.8% 74.7%
-Table 9: Quantitative complexity analysis for DFT Backbone with different layers and MLP Backbone+Projector for MindEye.
+$O(4nD^{\prime}\operatorname{log}n+(n+2nn^{\prime}+2n^{\prime})D^{\prime})$ (18)
+$=O((n\operatorname{log}n+nn^{\prime}+n^{\prime})D^{\prime})$
 Figure 13: T-SNE visualization between the voxel embeddings output by DFT Backbone and the image embedding of frozen CLIP.

### v3

@@
 Lite-Mind: Towards Efficient and Robust Brain Representation Learning
-Abstract
-Research in decoding visual information from the brain, particularly through the non-invasive fMRI method, is rapidly progressing.
-The challenge arises from the limited data availability and the low signal-to-noise ratio of fMRI signals, leading to a low-precision task of fMRI-to-image retrieval.
+Abstract.
+The limited data availability and the low signal-to-noise ratio of fMRI signals lead to the challenging task of fMRI-to-image retrieval.
 State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a large model, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s Vision Transformer (ViT).
-However, significant individual variations exist among subjects, even within identical experimental setups, mandating the training of large subject-specific models.
+However, significant individual variations exist among subjects, even under identical experimental setups, mandating the training of large subject-specific models.
 The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices.
-To this end, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm based on Discrete Fourier Transform, that efficiently aligns fMRI voxels to fine-grained information of CLIP.
+To this end, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm based on Discrete Fourier Transform (DFT), which efficiently aligns fMRI voxels to fine-grained information of CLIP.
+We elaborately design a DFT backbone with Spectrum Compression and Frequency Projector modules to learn informative and robust voxel embeddings.
 Our experiments demonstrate that Lite-Mind achieves an impressive 94.6% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye.
 Lite-Mind is also proven to be able to be migrated to smaller fMRI datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset.
-Code will be available at https://github.com/gongzix/Lite-Mind.
-1 Introduction
-Brain decoding holds immense significance in elucidating intricate mental processes and serves as a pivotal technology for brain-computer interfaces [26, 14, 28].
+1. Introduction
+Brain decoding holds immense significance in elucidating intricate mental processes and serves as a pivotal technology for brain-computer interfaces (Naselaris et al., 2011; Kamitani and Tong, 2005; Palazzo et al., 2020).
 Among diverse brain decoding tasks, the decoding of visual information stands out as a paramount yet challenging endeavor, allowing us to unravel the complex mechanisms involved in visual processing, object recognition, scene understanding, and even higher-order cognitive functions.
-In the pursuit of decoding natural visual information, functional magnetic resonance imaging (fMRI) is widely employed as a non-invasive modality to decipher the perceptual and semantic information intricately encoded within the cerebral cortex [22].
+In the pursuit of decoding natural visual information, functional magnetic resonance imaging (fMRI) is widely employed as a non-invasive modality to decipher the perceptual and semantic information intricately encoded within the cerebral cortex (Linden, 2021).
 It draws considerable attention to fMRI-to-image retrieval/reconstruction tasks.
 This paper focuses primarily on tasks related to fMRI-to-image retrieval.
-Figure 1: The significance of lightweight brain decoding model. Subject-specific models are visually represented by clouds of distinct colors, and the cloud size correlates with model parameters.
+Figure 1. The significance of lightweight brain decoding model. Subject-specific models are visually represented by clouds of distinct colors, and the cloud size correlates with model parameters.
 The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images.
 However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models.
-To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models [29, 15, 9, 1, 13, 48].
-Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models [30] as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures [23, 25, 21, 35].
-Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader [21].
-Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip [23] and UniBrain [25] introduce the text annotations of images to assist in aligning fMRI voxels with image representations.
-Brain-diffuser [27] aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
-Among the CLIP-related models, MindEye [35] stands out by employing a large-scale (996M) MLP Backbone and Projector to map fMRI voxels and align their representations with the final hidden layer of CLIP using contrastive learning.
+To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models (Pereira et al., 2018; Kewenig et al., 2023; Emmanuele et al., 2021; Allen et al., 2022; Horikawa and Kamitani, 2017; Zou et al., 2022).
+Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models (Radford et al., 2021) as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures (Liu et al., 2023; Mai and Zhang, 2023; Lin et al., 2022; Scotti et al., 2023).
+Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader (Lin et al., 2022).
+Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip (Liu et al., 2023) and UniBrain (Mai and Zhang, 2023) introduce the text annotations of images to assist in aligning fMRI voxels with image representations.
+Brain-diffuser (Ozcelik and VanRullen, 2023) aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
+Among the CLIP-related models, MindEye (Scotti et al., 2023) stands out by employing a large-scale (996M) MLP Backbone and Projector to map fMRI voxels and align their representations with the final hidden layer of CLIP using contrastive learning.
 MindEye achieves remarkable breakthroughs, surpassing $90\%$ accuracy for both image and brain retrieval, outperforming previous state-of-the-art accuracies that were below $50\%$.
@@
 This can lead to model parameter redundancy and decelerated convergence speed.
-Recently, the Fourier Transform has gained attention in deep learning, demonstrating lightweight characteristics and improved efficiency in learning frequency domain patterns, due to its global perspective and energy concentration properties [47].
+Recently, the Fourier Transform has gained attention in deep learning, demonstrating lightweight characteristics and improved efficiency in learning frequency domain patterns, due to its global perspective and energy concentration properties (Yi et al., 2024).
 Additionally, the Fourier Transform naturally possesses the advantage of processing signals: it is more effective and efficient to filter out noise and maintain informative voxel signals in the frequency domain.
 It presents a promising alternative to large-scale MLPs for encoding fMRI signals.
-To tackle the trade-off between model capacity and scalability in fMRI-to-image retrieval, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm.
+In order to mitigate the noise effect of fMRI and obtain more robust latent representations, we designed global filter-based Filter Blocks for fMRI denoising and fMRI spectrum compression.
+In addition to the global view property of frequency domain filtering, which can block out the noise and thus obtain better robust representations, the frequency domain computation also brings a great efficiency improvement compared to the large MLP structure.
+Since the features of the compressed fMRI we obtain are more concentrated in the frequency domain, we also design a Frequency Projector based on MLP in the frequency domain.
+Combining all our computational paradigms in frequency domain, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm.
 We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the large MLP Backbone used in MindEye.
@@
 Our contributions are summarized as follows:
-  • We demonstrated that the Fourier Transform has the advantage of lightweight and efficient for fMRI representation learning in frequency domain.
-  • We proposed a new Brain Representation Network DFT Backbone based on Discrete Fourier Transform, as an efficient and robust fMRI representation network for fMRI-to-image retrieval.
-Its effectiveness, efficiency, and lightweight are demonstrated.
+  • We demonstrated that the Fourier Transform has the advantage of being lightweight and efficient for fMRI representation learning in the frequency domain.
+  • We proposed a novel DFT Backbone with elaborate Spectrum Compression and Frequency Projector modules, which is theoretically and empirically verified lightweight, efficient, and robust for brain representation.
   • We migrated Lite-Mind to various downstream tasks, and demonstrated its robustness. It achieves state-of-the-art zero-shot classification performance on the GOD dataset.
-2 Related Work
-2.1 Brain Visual Decoding
+2. Related Work
+2.1. Brain Visual Decoding
 The study of visual decoding using fMRI in the human brain has been a long-standing endeavor.
 In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation is how to align fMRI signals to the image representation space or intermediate space.
-Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes [38].
-Further research utilizes pre-trained VGG to enrich hierarchical image features [13, 36, 7].
+Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes (Takagi and Nishimoto, 2023).
+Further research utilizes pre-trained VGG to enrich hierarchical image features (Horikawa and Kamitani, 2017; Shen et al., 2019; Du et al., 2023).
 With the development of cross-modal tasks, the image-text pre-training model CLIP has been introduced into fMRI-image research.
-Mind-reader [21] is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow resnet-like model to align fMRI with the CLS embeddings output by the CLIP image encoder.
+Mind-reader (Lin et al., 2022) is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow Resnet-like model to align fMRI with the CLS embeddings output by the CLIP image encoder.
 Mind-Reader did not achieve good performance in fMRI-image retrieval when using InfoNCE loss.
-BrainClip [23] uses a VAE model for retrieval tasks while still using the CLS embeddings.
-Brain diffuser [27] introduced a fine-grained representation of the image of the last hidden layer in CLIP.
-But they used 257 different linear regression models to align each layer of the 257 hidden layers, resulting in the model being too cumbersome.
-Mind-Vis [5] conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values.
-MindEye [35] uses a large MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 [31] for the first time to narrow the disjointed vectors.
+BrainClip (Liu et al., 2023) uses a VAE model for retrieval tasks while still using the CLS embeddings.
+Brain diffuser (Ozcelik and VanRullen, 2023) introduced a fine-grained representation of the image of the last hidden layer in CLIP.
+However they used 257 different linear regression models to align each layer of the 257 hidden layers, resulting in the model being too cumbersome.
+Mind-Vis (Chen et al., 2023) conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values.
+MindEye (Scotti et al., 2023) uses a large MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 (Ramesh et al., 2022) for the first time to narrow the disjointed vectors.
 At the same time, it creatively proposed a mixed contrast loss MixCo as a data augmentation method to make the large model converge.
-The recent trend in brain fMRI decoding research tends to use larger models to achieve better performance on downstream tasks [42, 16], ignoring the privacy and deployment issues discussed above.
+The recent trend in brain fMRI decoding research tends to use larger models to achieve better performance on downstream tasks (Xia et al., 2024; Kneeland et al., 2023), ignoring the privacy and deployment issues discussed above.
 However, to the best of our knowledge, no previous research has focused on lightweight and efficient brain networks.
-2.2 Fourier Transform in Deep Learning
+Figure 2.
+Overview of our Lite-Mind.
+Figures (a) and (b) show the architecture of the MLP Backbone of MindEye and the DFT backbone and Retrieval Pipeline, respectively.
+fMRI voxels are inputted into DFT Backbone to obtain voxel embeddings.
+2.2. Fourier Transform in Deep Learning
 Fourier Transform plays a vital role in the area of digital signal processing.
-It has been introduced to deep learning for enhanced learning performance [11, 8, 45], and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis.
-In the field of computer vision, GFNet [32] utilizes Fast Fourier Transform(FFT) to convert images to the frequency domain and exchange global information between learnable filters.
+It has been introduced to deep learning for enhanced learning performance (Gueguen et al., 2018; Ehrlich and Davis, 2019; Yang and Soatto, 2020), and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis.
+In the field of computer vision, GFNet (Rao et al., 2021) utilizes Fast Fourier Transform(FFT) to convert images to the frequency domain and exchange global information between learnable filters.
 As a continuous global convolution independent of input resolution, Guibas et al.
-[12] design the Adaptive Fourier Neural Operator(AFNO) frame token mixing.
+(Guibas et al., 2021) design the Adaptive Fourier Neural Operator(AFNO) frame token mixing.
 Xu et al.
-[43] devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy.
+(Xu et al., 2020) devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy.
 As for natural language processing, Lee-Thorp et al.
-[20] use the Fourier Transform as a text token mixing mechanism on text classification.
-In research of time series forecast, Fourier Transform has deeper applications  [3, 18, 17, 44, 46, 19].
+(Lee-Thorp et al., 2021) use the Fourier Transform as a text token mixing mechanism for text classification.
+In research of time series forecast, Fourier Transform has deeper applications  (Cao et al., 2020; Lange et al., 2021; Koç and Koç, 2022; Yang and Hong, 2022).
 To increase the accuracy of multivariate time-series forecasting, Cao et al.
-[3] propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain.
+(Cao et al., 2020) propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain.
 Yang et al.
-[44] propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum.
+(Yang and Hong, 2022) propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum.
 Yi et al.
-[47] prove that frequency-domain MLP is more efficient than time-domain MLP.
+(Yi et al., 2024) prove that frequency-domain MLP is more efficient than time-domain MLP.
 Processing fMRI frequency domain information is more helpful for understanding the mechanism of collaborative operation within the human brain.
 However, there is currently no research on processing fMRI signals in the frequency domain for representation learning.
-Figure 2: Overview of our Lite-Mind.
-Figure (a) is the architecture of the MLP Backbone of MindEye.
-Figure (b) is the architecture of the DFT backbone and Retrieval Pipeline.
-fMRI voxels are inputted into DFT Backbone to obtain voxel embeddings for downstream tasks.
-3 Lite-Mind
-3.1 Overview
-As illustrated in Figure 2, Lite-Mind comprises two main components: 1) DFT Backbone consists of fMRI Spectrum Compression and fMRI Token Decoupling in Frequency-domain, mapping flattened voxels to an intermediate space.
+3. Lite-Mind
+3.1. Overview
+As illustrated in Figure 2, Lite-Mind comprises two main components: 1) DFT Backbone consists of fMRI Spectrum Compression and Frequency Projector, mapping flattened voxels to an intermediate space.
 2) Retrieval pipeline contains Diffusion Projector, Contrastive Learning, and Downstream tasks with the voxel embeddings from the DFT backbone.
-3.2 DFT Backbone
+3.2. DFT Backbone
 Sampling an fMRI-image pair ($x$,$y$) from dataset $D$, $x$ denotes the fMRI data and $y$ denotes the paired image.
-3.2.1 fMRI Spectrum Compression
-Patchify and Embedding The fMRI data $x$ is one-dimensional data composed of voxels.
-we firstly divide $x$ into $n$ non-overlapping patches $x=[x_{1},x_{2},...,x_{n}]$ using zero-padding, and adopt convolutions with positional encoding to obtain multiple tokens $t=[t_{1},t_{2},...,t_{n}]$.
+3.2.1. fMRI Spectrum Compression
+Patchify and Embedding The fMRI data $x$ after ROI preprocessing and flattening is a one-dimensional long vector composed of voxels.
+We conduct patchify firstly because it has been verified as simple and effective for spatial/temporal representation of long sequences or high-dimensional vectors.
+At first, we divide $x$ into $n$ non-overlapping patches $x=[x_{1},x_{2},...,x_{n}]$ using zero-padding, and adopt convolutions with positional encoding to obtain multiple tokens $t=[t_{1},t_{2},...,t_{n}]$.
 Filter Block. The spectrum of voxel tokens is obtained by 1D DFT as follows:
-$X[k]=F(t)=\sum_{i=1}^{n}t_{i}e^{-ki(2\pi/n)j}$ (1)
+(1) $X[k]=F(t)=\sum_{i=1}^{n}t_{i}e^{-ki(2\pi/n)j}$
 Where $X\in\mathbb{C}^{n\times d}$ is a complex tensor, $X[k]$ is the spectrum of $t$ at the frequency $2\pi k/n$, $F(\cdot)$ denotes the 1D DFT along the voxel token dimension.
 $i$ denotes the $i$-th token, and $j$ denotes unit imaginary.
-After obtaining the spectrum of voxel tokens, we introduce learnable filters for filtering processing.
-Voxel spatial features are effectively consolidated within each frequency element, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain.
-We use $\textbf{K}=[\textbf{k}_{1},\textbf{k}_{2},...,\textbf{k}_{M}]$ to represent the filter library, where $M$ is the number of filters in the filter library.
-Spectrum $\hat{X}$ after filtering and compression is obtained as follows:
-$\hat{X}=\sum_{m=1}^{M}\frac{1}{n}{|X|}^{2}\odot\textbf{k}_{m}cos(\frac{(2m-1)\pi}{2M})$ (2)
+Voxel spatial features are effectively consolidated within each element of the frequency spectrum of voxel tokens, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain.
+Accordingly, we introduce learnable filters $\textbf{K}=[\textbf{k}_{1},\textbf{k}_{2},...,\textbf{k}_{M}]$ to filter and compress the spectrum as follows, where $\mathbf{K}$ denotes the filter library and $M$ is the number of filters in the library.
+(2) $\hat{X}=\sum_{m=1}^{M}\frac{1}{n}{|X|}^{2}\odot\textbf{k}_{m}cos(\frac{(2m-1)\pi}{2M})$
 where $\odot$ is the element-wise multiplication, $|X|^{2}$ is the power spectrum of $X$.
 The $|X|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective.
-It also facilitates voxel information concentration and noise filtering.
-$cos((2m-1)\pi/2M)$ adopts the filterling form of Discrete Cosine Transform (DCT), which compacts better energy and can aggregate the more important information in fMRI voxels.
-Its combination of application with the channel filter library K allows for efficient frequency domain feature compression and noise denoising for fMRI.
-Finally, we employ Inverse Discrete Fourier Transform (IDFT), denoted by $F^{-1}(\cdot)$, to convert the spectrum back into the spatial domain:
-$\hat{t}\leftarrow F^{-1}(\hat{X})$ (3)
-3.2.2 fMRI Token Decoupling in Frequency-domain
-To align voxel tokens to image tokens, we decouple the filtered and compressed fMRI tokens again in frequency domain.
-Similarly, we follow Equation (2) to convert $\hat{t}$ to the frequency domain, then decouple tokens by a MLP-like model in frequency domain(FreMLP).
+$cos((2m-1)\pi/2M)$ adopts the filtering form of Discrete Cosine Transform (DCT), which compacts better energy and can aggregate the significant information in fMRI voxels.
+Its combination with the channel filter library K allows for efficient feature compression and noise denoising for fMRI from the frequency domain.
+Finally, we employ Inverse Discrete Fourier Transform (IDFT) $F^{-1}(\cdot)$ to convert the spectrum back into the spatial domain:
+(3) $\hat{t}\leftarrow F^{-1}(\hat{X})$
+3.2.2. Frequency Projector
+The fMRI spectrum after filtering has a more concentrated feature in the frequency domain compared to the spatial domain, which is caused by the energy concentration effect of the frequency domain calculation.
+We therefore designed a Frequency Projector in the frequency domain to align the image tokens, which has the added benefit of maintaining all computations of our DFT Backbone in the frequency domain, where lightness and efficiency are guaranteed.
+To align voxel tokens to image tokens, we mapped the filtered and compressed fMRI tokens by an MLP-like projector in the frequency domain(FreMLP).
+Similarly, we follow Equation (2) to convert $\hat{t}$ to the frequency domain.
 FreMLP.
 Following Equation (2), the complex number output of $\hat{t}$ after DFT is $\hat{X}\in\mathbb{C}^{n\times d}$.
-Given a complex number weight matrix $\mathcal{W}\in\mathbb{C}^{n\times n^{\prime}}$, and a complex number bias $\mathcal{B}\in\mathbb{C}^{n^{\prime}}$, then the frequency-domain MLPs for token decoupling can be formulated as:
-$X^{\prime}={\sigma(\hat{X}^{T}\mathcal{W}+\mathcal{B})}^{T}$ (4)
+Given a complex number weight matrix $\mathcal{W}\in\mathbb{C}^{n\times n^{\prime}}$, and a complex number bias $\mathcal{B}\in\mathbb{C}^{n^{\prime}}$, then the FreMLP can be formulated as:
+(4) $X^{\prime}={\sigma(\hat{X}^{T}\mathcal{W}+\mathcal{B})}^{T}$
 Where $X^{\prime}\in\mathbb{C}^{n^{\prime}\times d}$ is the final output, and $\sigma$ is the activation function.
 As both $\hat{X}$ and $\mathcal{W}$ are complex numbers, according to the rule of multiplication of complex numbers(see Appendix D.1 for details), we further extend the Equation (5) to:
-$\begin{split}X^{\prime}&=({\sigma(Re(\hat{X}^{T})\mathcal{W}_{r}-Im(\hat{X}^{T})\mathcal{W}_{i}+\mathcal{B}_{r})}\\ &+j\sigma(Re(\hat{X}^{T})\mathcal{W}_{i}+Im(\hat{X}^{T})\mathcal{W}_{r}+\mathcal{B}_{i}))^{T}\end{split}$ (5)
+(5) $\begin{split}X^{\prime}&=({\sigma(Re(\hat{X}^{T})\mathcal{W}_{r}-Im(\hat{X}^{T})\mathcal{W}_{i}+\mathcal{B}_{r})}\\ &+j\sigma(Re(\hat{X}^{T})\mathcal{W}_{i}+Im(\hat{X}^{T})\mathcal{W}_{r}+\mathcal{B}_{i}))^{T}\end{split}$
 Where $Re(\cdot)$ denotes the real part of $\hat{X}^{T}$, $Im(\cdot)$ denotes the imaginary part, $\mathcal{W}=\mathcal{W}_{r}+j\mathcal{W}_{i}$ and $\mathcal{B}=\mathcal{B}_{r}+j\mathcal{B}_{i}$.
 Due to the more significant features of filtered and compressed tokens in the frequency domain, we replace the final FC layer of MindEye’s MLP Backbone with FreMLP.
-For more theoretical proof and implementation of FreMLP, please refer to the Appendix D.2.
+For more theoretical proof and implementation of FreMLP, please refer to Appendix D.1.
 Finally, we follow IDFT $t^{\prime}\leftarrow F^{-1}(X^{\prime})$ to get final tokens $t^{\prime}$, and use voxel embedding $f$ to represent them.
-3.3 Retrieval Pipeline
-3.3.1 Optimization Objective
-The optimization objective for the retrieval task is defined as follows:
+3.3. Retrieval Pipeline
+3.3.1. Optimization Objective
+The optimization objective for the retrieval task is as follows:
 […]
 where $\omega$ is the weight of our DFT backbone, $SIM(\cdot)$ denotes the cosine similarity, and $CLIP(\cdot)$ denotes using CLIP model to extract image embedding.
-3.3.2 Diffusion Projector
+3.3.2. Diffusion Projector
 When searching on a large-scale dataset like LAION-5B, image-to-image performs better than fMRI-to-image and can find more similar images.
-Therefore, we used a diffusion projector to translate the disjointed voxel CLS embeddings of the DFT Backbone output, using a DALLE · 2[31] similar diffusion model to obtain the image CLS embeddings and retrieve it on LAION-5B: $\mathcal{V^{\prime}}=Diffusion(f)$.
+Therefore, we used a diffusion projector to translate the disjointed voxel CLS embeddings of the DFT Backbone output, using a DALLE · 2(Ramesh et al., 2022) similar diffusion model to obtain the image CLS embeddings and retrieve it on LAION-5B: $\mathcal{V^{\prime}}=Diffusion(f)$.
 The Diffusion Projector consists of a transformer for image CLS embedding forecasting.
-3.3.3 Contrastive Learning
+3.3.3. Contrastive Learning
 Due to the inherent differences between different modalities, we use CLIP’s contrastive loss to train the DFT backbone.
 Sampling a batch $B$ from voxel-image pairs, the contrastive loss is defined below:
-$L_{contr}=-\frac{1}{|B|}\sum_{s=1}^{|B|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|B|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$ (7)
+(7) $L_{contr}=-\frac{1}{|B|}\sum_{s=1}^{|B|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|B|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$
 where $f_{s}$ is the $s$-th voxel embeddings, $\mathcal{V}_{s}$ is its corresponding image embeddings, $\tau$ is a temperature factor.
 To perform image-to-image retrieval on LAION-5B, we use MSE loss to constrain the generation of approximate image embeddings:
-$L_{mse}=\frac{1}{|B|}\sum_{s=1}^{|B|}{||\mathcal{V}_{s}-\mathcal{V^{\prime}}_{s}||}^{2}_{2}$ (8)
+(8) $L_{mse}=\frac{1}{|B|}\sum_{s=1}^{|B|}{||\mathcal{V}_{s}-\mathcal{V^{\prime}}_{s}||}^{2}_{2}$
 Accordingly, the training loss is defined below:
-$L=L_{contr}+\alpha L_{mse}$ (9)
+(9) $L=L_{contr}+\alpha L_{mse}$
 In general, when $\alpha=0$, Lite-Mind is trained to output voxel embeddings for retrieval on the NSD test set.
 However, for LAION-5B retrieval, $\alpha$ is non-zero to map the voxel CLS embeddings outputted by Lite-Mind to image CLS embeddings, enabling online retrieval.
-3.3.4 Downstream tasks
+3.3.4. Downstream tasks
 The retrieval process on various downstream tasks is shown in the upper right of Figure 2.
@@
 Image retrieval is performed on novel classes on the test set, and the similarities between the retrieved images and simple CLIP class text prompts are calculated for classification tasks.
-Figure 3: Partial retrieval results of Lite-Mind on all 982 test images for Subject 1.
+Figure 3.
+Partial retrieval results of Lite-Mind on all 982 test images for Subject 1.
 With 12.5M DFT Backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates.
 The number below each image represents the similarity score.
-See more cases include failure retrieval in Appendix C.4.
-4 Experiments
-4.1 Dataset
-Natural Scenes Dataset (NSD) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images from MSCOCO-2017 dataset [1], which contains images of complex natural scenes.
+See more cases including failure retrieval in Appendix C.4.
+4. Experiments
+4.1. Dataset
+Natural Scenes Dataset (NSD) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images from MSCOCO-2017 dataset (Allen et al., 2022), which contains images of complex natural scenes.
 Participants viewed three repetitions of 10,000 images with a 7-Tesla fMRI scanner over 30–40 sessions.
@@
 These regions span visual areas ranging from the early visual cortex to higher visual areas.
-Our experimental setup is consistent with the NSD image reconstruction and retrieval articles [38, 27, 25, 35, 21].
-Generic Object Decoding (GOD) Dataset [^2] http://brainliner.jp/data/brainliner/Generic_Object_Decoding was created by Horikawa and Kamitani [13], consisting of fMRI recordings of five healthy subjects who were presented with images from ImageNet [6].
+Our experimental setup is consistent with the NSD image reconstruction and retrieval articles (Takagi and Nishimoto, 2023; Ozcelik and VanRullen, 2023; Mai and Zhang, 2023; Scotti et al., 2023; Lin et al., 2022).
+Generic Object Decoding (GOD) Dataset [^2] http://brainliner.jp/data/brainliner/Generic_Object_Decoding was created by Horikawa and Kamitani (Horikawa and Kamitani, 2017), consisting of fMRI recordings of five healthy subjects who were presented with images from ImageNet (Deng et al., 2009).
 The GOD Dataset includes 1250 distinct images selected from 200 ImageNet categories.
@@
 For each test image, the fMRI responses from different trials were averaged.
-4.2 Implementation details
+4.2. Implementation details
 To fairly compare different backbones, we did not use any data augmentation methods, such as voxel mixing loss MixCo and image slicing enhancement.
@@
 More experimental details and hyperparameter settings can be found in Appendix A.
-5 Results
-5.1 fMRI/image retrieval
+Method Model Parameters Retrieval
+[…]
+Lin et al. (Lin et al., 2022) deep models 2.34M 11.0% 49.0%
+Ozcelik… (Ozcelik and VanRullen, 2023) 257 separate linear regression models 3B 29.9% 21.4%
+MindEye (Scotti et al., 2023) MLP Backbone 940M 89.6% 82.2%
+MindEye (Scotti et al., 2023) MLP Backbone+Projector 996M 88.8% 84.9%
+MindEye (Scotti et al., 2023) MLP Backbone+Prior 1.2B 93.4% 90.1%
+Lite-Mind(ours) DFT Backbone 12.5M 94.6% 97.1%
+(a) Quantitative comparison of Lite-Mind retrieval performance against other models for Subject 1.
+Image retrieval refers to the hit rate of correct retrieval from 300 candidates, given the associated brain sample (chance=0.3%); vice-versa for brain retrieval.
+Lite-Mind only uses a parameter quantity of 12.5M, achieving extremely high retrieval performance without using any model to close vectors of different modalities (see Appendix B.1 for remaining subject models).
+Method Low-Level High-Level
+[…]
+Lin et al. (Lin et al., 2022) - - - - 78.2% - - -
+Takagi… (Takagi and Nishimoto, 2023) - - 83.0% 83.0% 76.0% 77.0% - -
+Gu et al. (Gu et al., 2022) .150 .325 - - - - .862 .465
+Ozcelik… (Ozcelik and VanRullen, 2023) .254 .356 94.2% 96.2% 87.2% 91.5% .775 .423
+BrainCLIP (Liu et al., 2023) - - - - 86.7% 94.8% - -
+MindEye (Scotti et al., 2023) .309 .323 94.7% 97.8% 93.8% 94.1 .645 .367
+MindEye-125M(LAION-5B) .130 .308 84.0% 92.6% 86.9% 86.1% .778 .477
+Lite-Mind-8M(LAION-5B) .125 .331 78.7% 89.4% 87.9% 88.7% .724 .446
+(b) LAION-5B retrieval performance.
+The upper part is based on various fMRI-image reconstruction methods using generative neural networks, while the lower part represents the retrieval performance of the large-scale dataset LAION-5B as an alternative to fMRI-image reconstruction.
+All results are averaged across four subjects, and Lite-Mind achieved higher performance with a lighter backbone, even exceeding the accuracy of many reconstruction methods.
+Table 1. Main results of Lite-Mind retrieval and LAION-5B retrieval on the NSD dataset.
+5. Results
+5.1. fMRI/image retrieval
 Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings on the test set.
@@
 fMRI retrieval is the opposite process mentioned above.
-Note that there are many semantically and visually similar images on the NSD dataset test set, which are considered to be similar in CLIP space.
+Note that there are many semantically and visually similar images on the NSD test set, considered to be similar in CLIP space.
 Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image.
 For test retrieval, we adhered to the identical methodology as Lin et al.
-[21] and MindEye [35] to compute the retrieval metrics presented in Table 0(a).
+(Lin et al., 2022) and MindEye (Scotti et al., 2023) to compute the retrieval metrics presented in Table 1(a).
 For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images.
@@
 Note that our DFT Backbone is closest to MindEye’s MLP Backbone with only contrastive learning, while Prior in MLP Backbone+Prior includes MSE loss for image reconstruction.
-Method Model Parameters Retrieval
-[…]
-Lin et al. [21] deep models 2.34M 11.0% 49.0%
-Ozcelik… [27] 257 separate linear regression models 3B 29.9% 21.4%
-MindEye [35] MLP Backbone 940M 89.6% 82.2%
-MindEye [35] MLP Backbone+Projector 996M 88.8% 84.9%
-MindEye [35] MLP Backbone+Prior 1.2B 93.4% 90.1%
-Lite-Mind(ours) DFT Backbone 12.5M 94.6% 97.1%
-(a) Quantitative comparison of Lite-Mind retrieval performance against other models for Subject 1.
-Image retrieval refers to the hit rate of correct retrieval from 300 candidates, given the associated brain sample (chance=0.3%); vice-versa for brain retrieval.
-Lite-Mind only uses a parameter quantity of 12.5M, achieving extremely high retrieval performance without using any model to close vectors of different modalities(see Appendix B.1 for remaining subject models).
-Method Low-Level High-Level
-[…]
-Lin et al. [21] - - - - 78.2% - - -
-Takagi… [38] - - 83.0% 83.0% 76.0% 77.0% - -
-Gu et al. [10] .150 .325 - - - - .862 .465
-Ozcelik… [27] .254 .356 94.2% 96.2% 87.2% 91.5% .775 .423
-BrainCLIP [23] - - - - 86.7% 94.8% - -
-MindEye [35] .309 .323 94.7% 97.8% 93.8% 94.1 .645 .367
-MindEye-125M(LAION-5B) .130 .308 84.0% 92.6% 86.9% 86.1% .778 .477
-Lite-Mind-8M(LAION-5B) .125 .331 78.7% 89.4% 87.9% 88.7% .724 .446
-(b) LAION-5B retrieval performance.
-The upper part is based on various fMRI-image reconstruction methods using generative neural networks, while the lower part represents the retrieval performance of the large-scale dataset LAION-5B as an alternative to fMRI-image reconstruction.
-All results are averaged across four subjects, and Lite-Mind achieved higher performance with a lighter backbone, even exceeding the accuracy of many reconstruction methods.
-Table 1: Main retrieval results on the NSD and LAION-5B dataset.
-5.2 LAION-5B retrieval
-Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B [34].
-The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client [2].
+5.2. LAION-5B retrieval
+Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B (Schuhmann et al., 2022).
+The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client (Beaumont, 2022).
 For LAION-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS embeddings of CLIP.
@@
 The experimental results are shown in Table 1(b).
-The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric [40]; EfficientNetB1(”Eff”) [39] and SwAV-ResNet50(”SwAV”) [4] refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
-Experiments show that our DFT Backbone performs better than MindEye in high-level metrics, as shown in Table 0(b), although our backbone only has 8M parameters(MindEye-CLS backbone still has 125M.
+The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric (Wang et al., 2004); EfficientNetB1(”Eff”) (Tan and Quoc, [n.
+d.]) and SwAV-ResNet50(”SwAV”) (Caron et al., 2020) refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
+Figure 4.
+Retrieval results of Lite-Mind from LAION-5B for Subject 1.
+The left column marked by a red box in every two columns represents the original image seen by the Subject, and the right column represents the image retrieved from LAION-5B.
+Methods Modality Prompt Subject1 Subject2 Subject3 Subject4 Subject5
+top-1 top5 top-1 top5 top-1 top5 top-1 top5 top-1 top5
+CADA-VAE (Schonfeld et al., 2019) V&T - 6.31 35.70 6.45 40.12 17.74 54.34 12.17 36.64 7.45 35.04
+MVAE (Wu and Goodman, 2018) V&T - 5.77 31.51 5.40 38.46 17.11 52.46 14.02 40.90 7.89 34.63
+MMVAE (Shi et al., 2019) V&T - 6.63 38.74 6.60 41.03 22.11 56.28 14.54 42.45 8.53 38.14
+MoPoE-VAE (Loshchilov and Hutter, 2017) V&T - 8.54 44.05 8.34 48.11 22.68 61.82 14.57 58.51 10.45 46.40
+BraVL (Du et al., 2023) V&T - 9.11 46.80 8.91 48.86 24.00 62.06 15.08 60.00 12.86 47.94
+BrainClip-VAE (Liu et al., 2023) V&T Text 8.00 42.00 24.00 52.00 20.00 58.00 20.00 58.00 20.00 46.00
+BrainClip-VAE (Liu et al., 2023) V&T CoOp 14.00 54.00 20.00 60.00 21.33 64.67 16.67 66.67 18.00 52.00
+MindEye-211M (Scotti et al., 2023) V Text 6.00 40.00 20.00 60.00 28.00 70.00 14.00 54.00 10.00 64.00
+Lite-Mind-15.5M(ours) V Text 26.00 74.00 28.00 70.00 30.00 80.00 34.00 76.00 24.00 66.00
+Table 2.
+Zero-shot visual stimulus classification on the GOD dataset.
+The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%).
+The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from (Du et al., 2023).
+MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B.3 for retrieval performance on the GOD dataset).
+Experiments show that our DFT Backbone performs better than MindEye in high-level metrics, as shown in Table 1(b), although the backbone only has 8M parameters (MindEye-CLS backbone still has 125M.
 Note that both MindEye and Lite-Mind perform LAION-5B retrieval based on CLS alignment, resulting in a smaller number of parameters compared to full models).
 The results indicate that an efficient DFT backbone can also assist downstream retrieval tasks.
-Partial retrieval visualization results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval and visualization results of each Subject are shown in Appendix B.2.
-5.3 GOD zero-shot classification
-Since the collective fMRI data are much fewer and shorter than those of the NSD dataset usually, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify the Lite-Mind’s generalization ability.
-The voxel lengths collected by Subjects on the GOD dataset range from 4133 to 4643, approximately 30% of NSD, and there are only 1200 fMRI-image pairs on the training set (4.8% of NSD).
+The retrieval visualization results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval performance and visualization results of each Subject are shown in Appendix B.2.
+5.3. GOD zero-shot classification
+Since fMRI data are much fewer and shorter than those of the NSD dataset usually, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify Lite-Mind’s generalization ability.
+The voxel lengths collected by Subjects on the GOD dataset range from 4133 to 4643, approximately 30% of NSD, and there are only 1200 fMRI-image pairs on the training dataset (4.8% of NSD).
 We conduct zero-shot classification by using fMRI-to-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification).
@@
 On the contrary, due to the sudden decrease in dataset size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data size for voxel value-wise mappings with large MLP Backbone.
-More retrieval and classification results are in Appendix B.3.
-Figure 4: Partial retrieval results of Lite-Mind from LAION-5B for Subject 1.
-The left column marked by a red box in every two columns represents the original image seen by the Subject, and the right column represents the image retrieved from LAION-5B.
-Methods Modality Prompt Subject1 Subject2 Subject3 Subject4 Subject5
-top-1 top5 top-1 top5 top-1 top5 top-1 top5 top-1 top5
-CADA-VAE [33] V&T - 6.31 35.70 6.45 40.12 17.74 54.34 12.17 36.64 7.45 35.04
-MVAE [41] V&T - 5.77 31.51 5.40 38.46 17.11 52.46 14.02 40.90 7.89 34.63
-MMVAE [37] V&T - 6.63 38.74 6.60 41.03 22.11 56.28 14.54 42.45 8.53 38.14
-MoPoE-VAE [24] V&T - 8.54 44.05 8.34 48.11 22.68 61.82 14.57 58.51 10.45 46.40
-BraVL [7] V&T - 9.11 46.80 8.91 48.86 24.00 62.06 15.08 60.00 12.86 47.94
-BrainClip-VAE [23] V&T Text 8.00 42.00 24.00 52.00 20.00 58.00 20.00 58.00 20.00 46.00
-BrainClip-VAE [23] V&T CoOp 14.00 54.00 20.00 60.00 21.33 64.67 16.67 66.67 18.00 52.00
-MindEye-211M [35] V Text 6.00 40.00 20.00 60.00 28.00 70.00 14.00 54.00 10.00 64.00
-Lite-Mind-15.5M(ours) V Text 26.00 74.00 28.00 70.00 30.00 80.00 34.00 76.00 24.00 66.00
-Table 2: Zero-shot visual stimulus classification on the GOD dataset.
-The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%).
-The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from [7].
-MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B.3 for retrieval performance on the GOD dataset).
-5.4 Ablations and visualization
+More retrieval and classification results are provided in Appendix B.3.
+5.4. Ablations and visualization
 We assess the effectiveness of different modules within Lite-Mind to investigate the lightweight and efficient performance in this session. All experiments below are for Subject 1 on the NSD dataset.
@@
 We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of fMRI Spectrum Compression on retrieval efficiency(Table 3).
-As the Filter Blocks layers gradually deepen, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual.
+As the Filter Block layers gradually deepen, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual.
 We interestingly found that the accuracy of brain retrieval tends to converge faster than the image retrieval and only 6 layers of Filter Blocks are needed to approach the highest value.
@@
 21 0.80 12.5M 94.6% 97.4%
-Table 3: Ablation experiments on fMRI Spectrum Compression revealed a positive correlation between the depth of the Filter Blocks and enhanced forward retrieval accuracy.
+Table 3. Ablation experiments on fMRI Spectrum Compression revealed a positive correlation between the depth of the Filter Blocks and enhanced forward retrieval accuracy.
 Module Analysis.
 We conducted ablation experiments on two main modules of DFT Backbone in Table 4.
-We completely removed the Filter blocks and aligned tokens directly to the image embeddings with fMRI token decoupling, achieving a forward retrieval accuracy of 65.8%, indicating that the importance of frequency domain filtering for denoising and compression.
-Then we replaced FreMLP with the FC layer in the real domain, where tokens are flatten into a tensor for alignment, similar to the last layer of the MLP backbone of MindEye.
+We completely removed the Filter blocks and aligned tokens directly to the image embeddings with frequency projector, achieving a forward retrieval accuracy of 65.8%, indicating the importance of frequency domain filtering for denoising and compression.
+Then we replaced FreMLP with the FC layer in the real domain, where tokens are flattened into a tensor for alignment, similar to the last layer of the MLP backbone of MindEye.
 The model performance rapidly declined, indicating that tokens processed by DFT Backbone have better mapping performance with FreMLP in the frequency domain.
@@
 w/o FreMLP 43.4% 47.8%
-Table 4: Ablation experiments on modules of DFT Backbone.
+Table 4. Ablation experiments on modules of DFT Backbone.
 Embedding Dimension.We conducted additional experiments to verify the impact of embedding dimensions on retrieval accuracy, as shown in Table 5.
@@
 ViT-L/14 Hidden 257×768 12.5M 94.6% 97.4%
-Table 5: Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
+Table 5. Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
 Cerebral cortex.
-In order to explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s [38] method to extract the stream ROI on the NSD dataset, covering the visual cortex of nsdgenal, and trained the individual retrieval DFT Backbones for visual cortices (early, lateral, parietal, ventral).
+To explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s (Takagi and Nishimoto, 2023) method to extract the stream ROI on the NSD dataset, covering the visual cortex of nsdgenal, and trained the individual retrieval DFT Backbones for visual cortices (early, lateral, parietal, ventral).
 As shown in Table 6, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy.
-Although with voxel length of 5917, Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research.
-They demonstrated that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving fully connected MLP Backbones are unnecessary.
+Although with a voxel length of 5917, Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research.
+They demonstrated that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving that fully connected MLP Backbones are unnecessary.
 Interestingly, for other cortical regions, both lateral and ventral have some impacts on retrieval accuracy, while parietal shows little.
+Visualization.
+We visualize the CLS embeddings for LAION-5B using T-SNE in Figure 5.
+The figure shows that the diffusion projector successfully transformed the voxel CLS embeddings learned from contrastive learning into the image CLS embeddings, and the two kinds of embeddings are well fused for image-to-image retrieval on the LAION-5B dataset.
+We have also visualized the alignment process of voxel embeddings by contrastive learning on the NSD dataset in Appendix C.3 and learned filters of Filter Blocks in fMRI Spectrum Compression in Appendix C.5.
+Figure 5.
+T-SNE of embeddings of 982 test fMRI by diffusion projector on LAION-5B retrieval for Subject 1.
+In the figure, image-clip represents image-clip CLS embeddings, while voxel-fft and voxel-diffusion represent voxel CLS embeddings by DFT Backbone or Diffusion Projector.
+The diffusion projector plays a role in bringing vectors closer, enabling image-to-image retrieval on the LAION-5B dataset.
 Region Voxel Length Parameters Retrieval
@@
 ventral 7604 12.3M 18.1% 22.4%
-Table 6: Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
-Visualization.
-We visualize the CLS embeddings for LAION-5B using T-SNE in Figure 5.
-The figure shows that the diffusion projector successfully transformed the voxel CLS embeddings learned from contrastive learning into the image CLS embeddings, and the two kinds of embeddingss are well fused together for image-to-image retrieval on the LAION-5B dataset.
-We have also visualized the alignment process of voxel embeddings by contrastive learning on the NSD dataset in Appendix C.3, and learned filters of Filter Blocks in fMRI Spectrum Compression in Appendix C.5.
-Figure 5: T-SNE of embeddings of 982 test fMRI by diffusion projector on LAION-5B retrieval for Subject 1.
-In the figure, image-clip represents image CLIP CLS embeddings, while voxel-fft and voxel-diffusion represents voxel CLS embeddings by DFT Backbone or Diffusion Projector.
-The diffusion projector plays a role in bringing vectors closer, enabling image-to-image retrieval on the LAION-5B dataset.
-6 Limitations and Future Work
+Table 6. Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
+6. Limitations and Future Work
 Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models.
@@
 Further design beyond the model level is still needed for cross-Subject models, to achieve Subject-specific few-shot or even zero-shot.
-In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence.
+In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than on Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence.
 This provides a promising avenue for the interpretability of deep learning in neuroscience.
-7 Conclusion
+7. Conclusion
 In this work, we propose Lite-Mind, an extremely lightweight brain representation learning paradigm based on Fourier Transform for cross-modal fMRI-to-image retrieval.
@@
 Meanwhile, Lite-Mind’s dependence on dataset size and voxel length is less than larger models, demonstrating excellent robustness across different fMRI datasets.
-Supplementary Material
-Appendix A Additional Method Details
-A.1 Implementation details
-Data preprocessing.
-We downloaded the NSD dataset from the official website and used Takagi’s code to extract nsdgenal fMRI voxels, while Takagi extracted the stream region of the NSD dataset.
-We noticed that MindEye scaled the fMRI voxel values in advance, while we did not.
-The difference in fMRI voxels input data is shown in the example in Figure 6.
-NSD image files come from nsd-stimuli.hdf5 file and have a unified size of $425\times 425$.
-We did not perform any data augmentation on the image and straightly extracted the hidden layer representation (size of $257\times 768$) of the image through CLIP ViT-L/14 for training.
-Hyper-parameters.
-On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval.
-Owning to the lightweight nature of Lite-Mind, the batch size is set to 500, the learning rate is 1e-3, patch size is set to 480, and Filter library size is 4.
-Filter block layers are the same for all subjects.
-About LAION-5B retrieval, for DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, while the weight decay is 6.02e-2 for diffusion projector.
-CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, the batch size is set to 80 and the learning rate is 1e-3 for DFT Backbone while 1.1e-4 for diffusion projector.
-Note that hyper-parameters in Diffusion Projector are the value of the open-source DALLE·2 from https://github.com/lucidrains/DALLE2-pytorch.
-On the GOD dataset, the hyperparameters are the same as those on the NSD dataset, except for the batch size and patch size which are set to 1200 and 8 for all 5 subjects on the GOD dataset, respectively.
-(a) MindEye
-(b) Lite-Mind
-Figure 6: The fMRI averaged activities of MindEye and Lite-Mind responding to the same image, respectively. The figures only visualize the activities of the first 1000 voxels for illustration.
-Appendix B Additional Experiment Results
-B.1 Additional Results of Retrieval
-To demonstrate the applicability of our method, we also conducted experiments on three other Subjects, i.e., Subject 2, Subject 5, and Subject 7, on the NSD dataset, and the experimental results are shown in Table 7.
-In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths.
-It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90%, and the retrieval accuracies of Subjects 5 and 7 are also above 80%.
-The results prove that DFT Backbone can efficiently work on different subjects.
-Note that MindEye only includes the results of the overall model in Subjects 2, 5, and 7 in its Appendix, and it does not evaluate the effect of individual MLP Backbone.
-As a result, Table 7 does not present any evaluation results for MindEye.
-Method Voxel Length Parameters Retrieval
-[…]
-Lite-Mind(Subj 1) 15724 12.51M 94.6% 97.4%
-Lite-Mind(Subj 2) 14278 12.49M 94.1% 98.2%
-Lite-Mind(Subj 5) 13039 12.47M 80.5% 86.3%
-Lite-Mind(Subj 7) 12682 12.47M 81.7% 82.3%
-Table 7: Addtional retrieval performance for individual subjects on 982 test images of the NSD dataset.
-Method Low-Level High-Level
-[…]
-Lite-Mind(Subject 1) .134 .332 78.8% 88.9% 88.5% 88.8% .730 .451
-Lite-Mind(Subject 2) .120 .328 78.0% 89.4% 86.3% 87.4% .730 .446
-Lite-Mind(Subject 5) .123 .332 79.4% 90.0% 88.8% 89.9% .712 .440
-Lite-Mind(Subject 7) .121 .331 78.7% 88.8% 87.8% 88.5% .723 .448
-Table 8: LAION-5B retrieval alternative reconstruction performance for the specific subject.
-B.2 Additional Results of LAION-5B Retrieval
-In order to better reflect the retrieval performance of Lite-Mind on LAION-5B, we presented the performance indicators of LAION-5B retrieval substitution reconstruction for other subjects, i.e., Subject 2, Subject 5, and Subject 7, in Table 8, corresponding to the average performance of subjects in Table 0(b).
-Similarly, the visualization results of the other subjects in Figure 7 correspond to the image samples of Subject 1 in Figure 4.
-Based on the comprehensive table and graph, it can be found that Lite-Mind has good generalization on all four subjects, verifying the LAION-5B retrieval ability of Lite-Mind on different subjects.
-Meanwhile, as shown in Figure 7, the retrieval performance of LAION-5B completely depends on the retrieval accuracy of the CLS model in the test set, such that images retrieved incorrectly in the test set may also have retrieval bias on LAION-5B, for example treating a teddy bear as an image of a cat or dog as shown in Figure 7.
-However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS embedding models.
-In the future, it would be beneficial to explore models that improve the alignment of CLS embeddings or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
-Figure 7: Addtional retrieval results corresponding to Figure 4.
-The left column marked by a red box in every two columns represents the original image seen by the subject, and the right column represents the image retrieved on LAION-5B.
-B.3 Additional Results of Zero-shot Classification
-In the main body of the paper, we only demonstrated the zero-shot classification effect of Lite-Mind on the GOD dataset.
-The corresponding retrieval results of each Subject are shown in the Table 9 below.
-[…]
-top1 top5
-Lite-Mind(Subj 1) 4466 15.50M 30.0% 60.0%
-Lite-Mind(Subj 2) 4404 15.46M 38.0% 58.0%
-Lite-Mind(Subj 3) 4643 15.64M 38.0% 72.0%
-Lite-Mind(Subj 4) 4133 15.24M 42.0% 62.0%
-Lite-Mind(Subj 5) 4370 15.43M 26.0% 54.0%
-Table 9: Retrieval performance on the GOD dataset.
-B.4 Hyper-parameters Experiment
-In this section, we explore the influence of some hyperparameters on the model, including patch size and number of filters $M$, to verify the parameter sensitivity of the model.
-All the experimental results on the NSD dataset are from Subject 1, with a retrieving pool size of 300.
-Patch size.
-We conducted experiments by varying the patch size as presented in Table 10.
-The results exhibit stable retrieval accuracy above 90%, which indicates that our DFT Backbone is not sensitive to part size.
-Since we found patch size has less effect on retrieval accuracy, we chose a relatively large patch size (e.g., 480) to ensure fewer parameters and faster convergence.
-Patch size Parameters Retrieval
-[…]
-50 14.0M 91.6% 96.4%
-200 12.7M 92.9% 96.9%
-480 12.5M 94.6% 97.1%
-600 12.3M 94.2% 97.0%
-900 11.9M 90.1% 95.5%
-Table 10: Retrieval performance for different patch size on the NSD dataset.
-Filter library size.
-We conducted experiments by varying the Filter library size as presented in Table 11.
-Specifically, there is a significant performance improvement from $M=1$ to $M=4$, while a slight fluctuation is observed for $M=4/8$.
-By setting $M=4$, the model has the ability to acquire diverse and distinct feature patterns from various dimensions of the frequency response while still maintaining an appropriate computational cost.
-Therefore, we determine that $M=4$ is the optimal choice on the NSD dataset.
-Filters Parameters Retrieval
-[…]
-1 9.0M 89.8% 96.4%
-2 10.2M 93.5% 97.4%
-4 12.5M 94.6% 97.1%
-8 17.2M 94.4% 96.7%
-Table 11: Retrieval performance for different Filter library size on the NSD dataset.
-Figure 8: Lite-Mind’s training loss curve and testing accuracy curve for Suject 1, 2, 5, 7 on the NSD dataset. The testing accuracy is calculated based on a retrieval pool of size 982.
-Figure 9: Lite-Mind’s retrieval heatmap on the NSD dataset for Subject 1.
-The larger figure on the right represents 982 × 982’s heatmap, and the smaller figure on the left represents the 50 × 50 subgraph.
-Appendix C Additional Visualization
-C.1 Training Curve
-We visualized the training process of all four subjects on the NSD dataset in Figure 8.
-As the training epochs increased, the loss of the training set rapidly decreased, while the accuracy of the test set rapidly increased and then showed a slow upward trend.
-Accuracy refers to the hit rate of correct retrieval from 982 test set images.
-C.2 Retrieval Heatmap
-We visualized the retrieval heatmap for Subject 1 on all 982 test images of the NSD dataset in Figure 9.
-It can be observed that the similarity is highest on the diagonal, and the color of the retrieved heat map is darker.
-It shows that Lite-Mind has effectively retrieved corresponding images, even if there are many similar images in the test set, which verifies the fine-grained ability of Lite-Mind.
-C.3 Information Alignment
-We visualized the T-SNE plot between the voxel embeddings output by DFT Backbone and the image embeddings of frozen CLIP as the accuracy of the test set improved, as shown in Figure 13.
-We can observe that as the training progresses, the retrieval accuracy of the test set improves, and the shape of voxel embeddings tends to be closer to image embeddings, indicating the success of contrastive learning.
-C.4 More Retrieval Cases
-In this section, we visualize retrieval failure cases in Figure 10, although only dozens of images are not in the Top 1.
-From the left half of the Figure, it can be seen that though not in Top 1, Lite-Mind still retrieved ground-truth in Top 2, and images in Top 4 are similar(either semantically similar or structurally similar, eg.
-animals in the wild or an airplane on the runway).
-On the contrary, a smaller proportion of ground-truth did not appear within the Top 4, as shown in the right half of the Figure.
-From the perspective of the images themselves, most of the scenes are too complex, even abstract (as shown in the second image), which may be the reason for the retrieval failure.
-Figure 10: Partial failure retrieval results of Lite-Mind on all 982 test images for Subject 1. The number below each image represents the similarity score.
-(a) The real part and imaginary of a filter weights.
-(b) The real part and imaginary of another filter weights from the same Filter Library.
-Figure 11: Weights visualization for DFT Backbone of $257\times 768$ embedding length.
-(a) The real part and imaginary of a filter weights.
-(b) The real part and imaginary of another filter weights from the same Filter Library.
-Figure 12: Weights visualization for DFT Backbone of 768 CLS embedding length.
-C.5 Visualization in Frequency Domain
-We visualized the weights of Filter library in Filter Blocks, as shown in Figure 11 and 12.
-Visualization is divided into the real part (left) and imaginary part (right) of filter weights.
-It can be observed that different filters have varying degrees of attention to different tokens, and the frequency domain better captures this characteristic.
-Interestingly, the weight of the imaginary part for the first and last tokens is almost always 0, indicating that the noise is distributed in these two tokens.
-Appendix D Theoretical Analysis.
-D.1 Complex Multiplication.
-For two complex number values $\mathcal{Z}_{1}=(a+jb)$ and $\mathcal{Z}_{2}=(c+jd)$, where $a$ and $c$ is the real part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively, $b$ and $d$ is the imaginary part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively.
-Then the multiplication of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ is calculated by:
-$\mathcal{Z}_{1}\mathcal{Z}_{2}$ $=(a+jb)(c+jd)$ (10)
-$=ac+j^{2}bd+jad+jbc$
-$=(ac-bd)+j(ad+bc)$
-D.2 Theorem Proof.
-Theorem 1.
-Suppose that H is the representation of raw fMRI voxel tokens and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of voxel tokens in the spatial domain is equal to the energy of its representation in the frequency domain.
-Formally, we can express this with the above notations:
-$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$ (11)
-Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the token dimension, $f$ is the frequency dimension.
-Proof.
-Given the representation of raw voxel token series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
-$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{\textbf{H}(v)}{\textbf{H}^{*}(v)}\,\mathrm{d}v$ (12)
-where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}\textbf{}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
-$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v&=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v\\ &=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f\\ &=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f\\ &=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f\end{aligned}$
-Proved. ∎
-Therefore, the energy of a voxel token series in the spatial domain is equal to the energy of its representation in the frequency domain.
-Theorem 2.
-Given the voxel token series input H and its corresponding frequency domain conversion H, the operations of frequency-domain MLP on H can be represented as global convolutions on H in the spatial domain.
-This can be given by:
-$\mathcal{H}\mathcal{W}+\mathcal{B}=\mathcal{F}(\textbf{H}*W+B)$ (14)
-Where $\mathcal{F}$ is DFT, *** is a circular convolution, $\mathcal{W}$ and $\mathcal{B}$ are the complex number weight and bias, and $W$ and $B$ are the weight and bias in the spatial domain.
-Proof.
-Suppose that we conduct operations in the N (i.e., token dimension), then
-$\mathcal{F}(\textbf{H}(v)*W(v))=\int_{-\infty}^{\infty}{(\textbf{H}(v)*W(v))e^{-j2\pi fv}}\,\mathrm{d}v$ (15)
-According to convolution theorem, $\textbf{H}(v)*W(v)=\int_{-\infty}^{\infty}(\textbf{H}(\tau)W(v-\tau))\,\mathrm{d}\tau$, then
-$\mathcal{F}(\textbf{H}(v)*W(v))&=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{(\textbf{H}(\tau)W(v-\tau))e^{-j2\pi fv}}\,\mathrm{d}\tau}\,\mathrm{d}v\\ &=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(v-\tau)e^{-j2\pi fv}}\,\mathrm{d}v}\textbf{H}(\tau)\,\mathrm{d}\tau\end{aligned}$
-Let $x=v-\tau$,then
-$\mathcal{F}(\textbf{H}(v)*W(v))&=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi f(x+\tau)}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi fx}e^{-j2\pi f\tau}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{\infty}{(\textbf{H}(\tau)*)e^{-j2\pi f\tau}}\,\mathrm{d}\tau\int_{-\infty}^{\infty}{(\textbf{W}(x)*)e^{-j2\pi fx}}\,\mathrm{d}x\\ &=\mathcal{H}(f)\mathcal{W}(f)\end{aligned}$
-Accordingly, $\textbf{H}(v)*W(v)$ in the spatial domain is equal to $\mathcal{H}(f)\mathcal{W}(f)$ in the frequency domain.
-Therefore, the operations of FreMLP ($\mathcal{H}\mathcal{W}+\mathcal{B}$) in the token dimension (i.e., v = N ) are equal to the operations ($\mathcal{F}(\textbf{H}*W+B)$) in the spatial domain.
-This implies that frequency-domain MLPs can be viewed as global convolutions in the spatial domain.
-Proved.
-∎
-D.3 Complexity Analysis
-For a fMRI voxel with a length of $l$, we divide it into $n$ patches.
-Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP Backbone and DFT Backbone respectively, the middle layer dimension of MLP Backbone is $D$, and the alignment embedding dimension is $n^{\prime}\times D^{\prime}$, where $n^{\prime}$ is the the number of tokens of CLIP.
-The time complexity of MLP Backbone is $O(lD+L_{1}D^{2}+n^{\prime}DD^{\prime})$.
-For DFT Backobone, the time complexity of patchify and tokenization is $O(lD^{\prime})$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nD^{\prime}\operatorname{log}n+nD^{\prime})$.
-The time complexity of FreMLP is $O(2nD^{\prime}\operatorname{log}n+2nn^{\prime}D^{\prime}+2n^{\prime}D^{\prime})$.
-Thus the time complexity of the entire DFT Backbone is:
-$O(4nD^{\prime}\operatorname{log}n+(n+2nn^{\prime}+2n^{\prime})D^{\prime})$ (18)
-$=O((n\operatorname{log}n+nn^{\prime}+n^{\prime})D^{\prime})$
-Figure 13: T-SNE visualization between the voxel embeddings output by DFT Backbone and the image embedding of frozen CLIP.
-Accuracy in the title refers to the hit rate of correct retrieval from 982 test set images and the blue dots represent voxel embeddings.

### v4

@@
 Lite-Mind is also proven to be able to be migrated to smaller fMRI datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset.
+fMRI, Brain-computer Interface (BCI), Cross-modal Retrieval
+copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681229isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
 1. Introduction
@@
 Figure 1. The significance of lightweight brain decoding model. Subject-specific models are visually represented by clouds of distinct colors, and the cloud size correlates with model parameters.
-The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images.
+The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images, following a series of cross-modal alignment efforts (Ramesh et al., 2022; Wu et al., 2024; Radford et al., 2021).
 However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models.
 To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models (Pereira et al., 2018; Kewenig et al., 2023; Emmanuele et al., 2021; Allen et al., 2022; Horikawa and Kamitani, 2017; Zou et al., 2022).
-Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models (Radford et al., 2021) as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures (Liu et al., 2023; Mai and Zhang, 2023; Lin et al., 2022; Scotti et al., 2023).
+Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models (Radford et al., 2021) as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures (Liu et al., 2023; Mai and Zhang, 2023; Lin et al., 2022; Scotti et al., 2023; Zhang et al., 2024).
 Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader (Lin et al., 2022).
-Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip (Liu et al., 2023) and UniBrain (Mai and Zhang, 2023) introduce the text annotations of images to assist in aligning fMRI voxels with image representations.
+BrainClip (Liu et al., 2023) and UniBrain (Mai and Zhang, 2023) introduce the text annotations of images to assist in aligning fMRI voxels with image representations.
 Brain-diffuser (Ozcelik and VanRullen, 2023) aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@
 Fourier Transform plays a vital role in the area of digital signal processing.
-It has been introduced to deep learning for enhanced learning performance (Gueguen et al., 2018; Ehrlich and Davis, 2019; Yang and Soatto, 2020), and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis.
+It has been introduced to deep learning for enhanced learning performance (Yi et al., 2023a; Lao et al., 2024; Yang and Soatto, 2020; Zhang et al., 2024), and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis.
 In the field of computer vision, GFNet (Rao et al., 2021) utilizes Fast Fourier Transform (FFT) to convert images to the frequency domain and exchange global information between learnable filters.
@@
 (Lee-Thorp et al., 2021) use the Fourier Transform as a text token mixing mechanism for text classification.
-In research of time series forecast, Fourier Transform has deeper applications  (Cao et al., 2020; Lange et al., 2021; Koç and Koç, 2022; Yang and Hong, 2022).
+In research of time series forecast, Fourier Transform has deeper applications  (Yi et al., 2023b; Lange et al., 2021; Koç and Koç, 2022).
 To increase the accuracy of multivariate time-series forecasting, Cao et al.
@@
 Finally, we follow IDFT $t^{\prime}\leftarrow F^{-1}(X^{\prime})$ to get final tokens $t^{\prime}$, and use voxel embedding $f$ to represent them.
+Figure 3.
+Partial retrieval results of Lite-Mind on all 982 test images for Subject 1.
+With 12.5M DFT Backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates.
+The number below each image represents the similarity score.
+See more cases including failure retrieval in Appendix C.4.
 3.3. Retrieval Pipeline
@@
 Image retrieval is performed on novel classes on the test set, and the similarities between the retrieved images and simple CLIP class text prompts are calculated for classification tasks.
-Figure 3.
-Partial retrieval results of Lite-Mind on all 982 test images for Subject 1.
-With 12.5M DFT Backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates.
-The number below each image represents the similarity score.
-See more cases including failure retrieval in Appendix C.4.
 4. Experiments
@@
 The detailed results and discussions are in Appendix B.2.
-As shown in Table 1(a), compared to MLP Backbone, our DFT Backbone improves retrieval accuracy by 5% and 14.9% for two retrieval ways, indicating that the fine-grained representation of the image comes from the rich representation of the last hidden layer of CLIP rather than MLP’s excessive attention to each voxel value.
-Note that our DFT Backbone is closest to MindEye’s MLP Backbone with only contrastive learning, while Prior in MLP Backbone+Prior includes MSE loss for image reconstruction.
-5.2. LAION-5B retrieval
-Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B (Schuhmann et al., 2022).
-The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client (Beaumont, 2022).
-For LAION-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS embeddings of CLIP.
-For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS embeddings, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP).
-Unlike MindEye, in the LAION-5B retrieval process, we used the Diffusion Projector to conduct image-to-image retrieval.
-To compare the performance of large-scale dataset retrieval, we use metrics of image reconstruction for evaluation with previous image reconstruction methods.
-The experimental results are shown in Table 1(b).
-The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric (Wang et al., 2004); EfficientNetB1(”Eff”) (Tan and Quoc, [n.
-d.]) and SwAV-ResNet50(”SwAV”) (Caron et al., 2020) refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
 Figure 4.
@@
 MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B.3 for retrieval performance on the GOD dataset).
+As shown in Table 1(a), compared to MLP Backbone, our DFT Backbone improves retrieval accuracy by 5% and 14.9% for two retrieval ways, indicating that the fine-grained representation of the image comes from the rich representation of the last hidden layer of CLIP rather than MLP’s excessive attention to each voxel value.
+Note that our DFT Backbone is closest to MindEye’s MLP Backbone with only contrastive learning, while Prior in MLP Backbone+Prior includes MSE loss for image reconstruction.
+5.2. LAION-5B retrieval
+Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B (Schuhmann et al., 2022).
+The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client (Beaumont, 2022).
+For LAION-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS embeddings of CLIP.
+For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS embeddings, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP).
+Unlike MindEye, in the LAION-5B retrieval process, we used the Diffusion Projector to conduct image-to-image retrieval.
+To compare the performance of large-scale dataset retrieval, we use metrics of image reconstruction for evaluation with previous image reconstruction methods.
+The experimental results are shown in Table 1(b).
+The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric (Wang et al., 2004); EfficientNetB1(”Eff”) (Tan and Quoc, [n.
+d.]) and SwAV-ResNet50(”SwAV”) (Caron et al., 2020) refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
 Experiments show that our DFT Backbone performs better than MindEye in high-level metrics, as shown in Table 1(b), although the backbone only has 8M parameters (MindEye-CLS backbone still has 125M.
@@
 Table 4. Ablation experiments on modules of DFT Backbone.
-Embedding Dimension.We conducted additional experiments to verify the impact of embedding dimensions on retrieval accuracy, as shown in Table 5.
-Among them, two different CLIP models (i.e., CLIP ViT/B-32 and CLIP ViT/L-14) were used to extract image embeddings, with a total of four embedding dimensions, by CLS embeddings and the last hidden layer, respectively.
-From the results, it can be seen that the retrieval accuracy of Lite-Mind is higher when the embedding dimensions of the image are longer and the representation of the image is richer.
-The results verify that the fine-grained alignment we mentioned above comes from the rich representation of the image, rather than the fMRI voxel value fully connected heavy MLP Backbone.
-Embeddings Dimension Parameters Retrieval
-[…]
-ViT-B/32 CLS 512 6.7M 57.5% 61.8%
-ViT-L/14 CLS 768 6.7M 60.3% 64.2%
-ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
-ViT-L/14 Hidden 257×768 12.5M 94.6% 97.4%
-Table 5. Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
 Cerebral cortex.
@@
 Table 5. Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
-6. Limitations and Future Work
+6. Limitations
 Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models.
 Although our exploratory experiments have shown that a common model can achieve comparable retrieval accuracy on four Subjects, it’s still necessary to ensure sufficient training data.
-Further design beyond the model level is still needed for cross-Subject models, to achieve Subject-specific few-shot or even zero-shot.
-In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than on Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence.
-This provides a promising avenue for the interpretability of deep learning in neuroscience.
 7. Conclusion
-In this work, we propose Lite-Mind, an extremely lightweight brain representation learning paradigm based on Fourier Transform for cross-modal fMRI-to-image retrieval.
+In this work, we propose Lite-Mind, an extremely lightweight brain representation learning paradigm based on Fourier Transform for fMRI-to-image retrieval.
 Our DFT Backbone is an efficient means of obtaining fine-grained representation alignment between fMRI signals and visual stimuli.
-With high retrieval accuracy, Lite-Mind can transform many downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification.
-Meanwhile, Lite-Mind’s dependence on dataset size and voxel length is less than larger models, demonstrating excellent robustness across different fMRI datasets.
+With high retrieval accuracy, Lite-Mind can transform downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification.
+Meanwhile, Lite-Mind’s dependence on dataset size and voxel length is less than that of larger models, demonstrating excellent robustness across different fMRI datasets.
+8. Acknowledgments
+This work is supported by the National Key Research and Development Program of China (No.
+2022YFB3104700), the National Natural Science Foundation of China (No.
+61976158, No.
+62376198), Shanghai Baiyulan Pujiang Project (No.
+08002360429).
+Appendix A Additional Method Details
+A.1. Implementation details
+Data preprocessing.
+We downloaded the NSD dataset from the official website and used Takagi’s code to extract nsdgenal fMRI voxels, while Takagi extracted the stream region of the NSD dataset.
+We noticed that MindEye scaled the fMRI voxel values in advance, while we did not.
+The difference in fMRI voxels input data is shown in the example in Figure 6.
+NSD image files come from nsd-stimuli.hdf5 file and have a unified size of $425\times 425$.
+We did not perform any data augmentation on the image and straightly extracted the hidden layer representation (size of $257\times 768$) of the image through CLIP ViT-L/14 for training.
+Hyper-parameters.
+On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval.
+Owning to the lightweight nature of Lite-Mind, the batch size is set to 500, the learning rate is 1e-3, patch size is set to 480, and Filter library size is 4.
+Filter block layers are the same for all subjects.
+About LAION-5B retrieval, for DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, while the weight decay is 6.02e-2 for diffusion projector.
+CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, the batch size is set to 80 and the learning rate is 1.16e-3 for DFT Backbone while 1.1e-4 for diffusion projector.
+Note that hyper-parameters in Diffusion Projector are the value of the open-source DALLE·2 from https://github.com/lucidrains/DALLE2-pytorch.
+On the GOD dataset, the hyperparameters are the same as those on the NSD dataset, except for the batch size and patch size which are set to 1200 and 8 for all 5 subjects on the GOD dataset, respectively.
+(a) MindEye
+(b) Lite-Mind
+Figure 6. The fMRI averaged activities of MindEye and Lite-Mind responding to the same image, respectively. The figures only visualize the activities of the first 1000 voxels for illustration.
+Appendix B Additional Experiment Results
+B.1. Additional Results of Retrieval
+To demonstrate the applicability of our method, we also conducted experiments on three other Subjects, i.e., Subject 2, Subject 5, and Subject 7, on the NSD dataset, and the experimental results are shown in Table 7.
+In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths.
+It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90%, and the retrieval accuracies of Subjects 5 and 7 are also above 80%.
+The results prove that DFT Backbone can efficiently work on different subjects.
+Note that MindEye only includes the results of the overall model in Subjects 2, 5, and 7 in its Appendix, and it does not evaluate the effect of individual MLP Backbone.
+As a result, Table 7 does not present any evaluation results for MindEye.
+Method Voxel Length Parameters Retrieval
+[…]
+Lite-Mind(Subj 1) 15724 12.51M 94.6% 97.4%
+Lite-Mind(Subj 2) 14278 12.49M 94.1% 98.2%
+Lite-Mind(Subj 5) 13039 12.47M 80.5% 86.3%
+Lite-Mind(Subj 7) 12682 12.47M 81.7% 82.3%
+Table 7. Addtional retrieval performance for individual subjects on 982 test images of the NSD dataset.
+Method Low-Level High-Level
+[…]
+Lite-Mind(Subject 1) .134 .332 78.8% 88.9% 88.5% 88.8% .730 .451
+Lite-Mind(Subject 2) .120 .328 78.0% 89.4% 86.3% 87.4% .730 .446
+Lite-Mind(Subject 5) .123 .332 79.4% 90.0% 88.8% 89.9% .712 .440
+Lite-Mind(Subject 7) .121 .331 78.7% 88.8% 87.8% 88.5% .723 .448
+Table 8. LAION-5B retrieval alternative reconstruction performance for the specific subject.
+B.2. Additional Results of LAION-5B Retrieval
+In order to better reflect the retrieval performance of Lite-Mind on LAION-5B, we presented the performance indicators of LAION-5B retrieval substitution reconstruction for other subjects, i.e., Subject 2, Subject 5, and Subject 7, in Table 8, corresponding to the average performance of subjects in Table 1.
+Similarly, the visualization results of the other subjects in Figure 7 correspond to the image samples of Subject 1 in Figure 4.
+Based on the comprehensive table and graph, it can be found that Lite-Mind has good generalization on all four subjects, verifying the LAION-5B retrieval ability of Lite-Mind on different subjects.
+Meanwhile, as shown in Figure 7, the retrieval performance of LAION-5B completely depends on the retrieval accuracy of the CLS model in the test set, such that images retrieved incorrectly in the test set may also have retrieval bias on LAION-5B, for example treating a teddy bear as an image of a cat or dog as shown in Figure 7.
+However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS embedding models.
+In the future, it would be beneficial to explore models that improve the alignment of CLS embeddings or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
+Figure 7.
+Addtional retrieval results corresponding to Figure 4.
+The left column marked by a red box in every two columns represents the original image seen by the subject, and the right column represents the image retrieved on LAION-5B.
+B.3. Additional Results of Zero-shot Classification
+In the main body of the paper, we only demonstrated the zero-shot classification effect of Lite-Mind on the GOD dataset.
+The corresponding retrieval results of each Subject are shown in the Table 9 below.
+[…]
+top1 top5
+Lite-Mind(Subj 1) 4466 15.50M 30.0% 60.0%
+Lite-Mind(Subj 2) 4404 15.46M 38.0% 58.0%
+Lite-Mind(Subj 3) 4643 15.64M 38.0% 72.0%
+Lite-Mind(Subj 4) 4133 15.24M 42.0% 62.0%
+Lite-Mind(Subj 5) 4370 15.43M 26.0% 54.0%
+Table 9. Retrieval performance on the GOD dataset.
+B.4. Hyper-parameters Experiment
+In this section, we explore the influence of some hyperparameters on the model, including patch size and number of filters $M$, to verify the parameter sensitivity of the model.
+All the experimental results on the NSD dataset are from Subject 1, with a retrieving pool size of 300.
+Patch size.
+We conducted experiments by varying the patch size as presented in Table 10.
+The results exhibit stable retrieval accuracy above 90%, which indicates that our DFT Backbone is not sensitive to part size.
+Since we found patch size has less effect on retrieval accuracy, we chose a relatively large patch size (i.e., 480) to ensure fewer parameters and faster convergence.
+Patch size Parameters Retrieval
+[…]
+50 14.0M 91.6% 96.4%
+200 12.7M 92.9% 96.9%
+480 12.5M 94.6% 97.1%
+600 12.3M 94.2% 97.3%
+900 11.9M 92.1% 95.5%
+Table 10. Retrieval performance for different patch size on the NSD dataset.
+Filter library size.
+We conducted experiments by varying the Filter library size as presented in Table 11.
+Specifically, there is a significant performance improvement from $M=1$ to $M=4$, while a slight fluctuation is observed for $M=4/8$.
+By setting $M=4$, the model has the ability to acquire diverse and distinct feature patterns from various dimensions of the frequency response while still maintaining an appropriate computational cost.
+Therefore, we determine that $M=4$ is the optimal choice on the NSD dataset.
+Filters Parameters Retrieval
+[…]
+1 9.0M 89.8% 96.4%
+2 10.2M 93.5% 97.4%
+4 12.5M 94.6% 97.1%
+8 17.2M 94.4% 96.7%
+Table 11. Retrieval performance for different Filter library size on the NSD dataset.
+Embedding Dimension.We conducted additional experiments to verify the impact of embedding dimensions on retrieval accuracy, as shown in Table 12.
+Among them, two different CLIP models (i.e., CLIP ViT/B-32 and CLIP ViT/L-14) were used to extract image embeddings, with a total of four embedding dimensions, by CLS embeddings and the last hidden layer, respectively.
+From the results, it can be seen that the retrieval accuracy of Lite-Mind is higher when the embedding dimensions of the image are longer and the representation of the image is richer.
+The results verify that the fine-grained alignment we mentioned above comes from the rich representation of the image, rather than the fMRI voxel value fully connected heavy MLP Backbone.
+Embeddings Dimension Parameters Retrieval
+[…]
+ViT-B/32 CLS 512 6.7M 57.5% 61.8%
+ViT-L/14 CLS 768 6.7M 60.3% 64.2%
+ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
+ViT-L/14 Hidden 257×768 12.5M 94.6% 97.4%
+Table 12. Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
+Figure 8. Lite-Mind’s training loss curve and testing accuracy curve for Suject 1, 2, 5, 7 on the NSD dataset. The testing accuracy is calculated based on a retrieval pool of size 982.
+Figure 9.
+Lite-Mind’s retrieval heatmap on the NSD dataset for Subject 1.
+The larger figure on the right represents 982 × 982’s heatmap, and the smaller figure on the left represents the 50 × 50 subgraph.
+Appendix C Additional Visualization
+C.1. Training Curve
+We visualized the training process of all four subjects on the NSD dataset in Figure 8.
+As the training epochs increased, the loss of the training set rapidly decreased, while the accuracy of the test set rapidly increased and then showed a slow upward trend.
+Accuracy refers to the hit rate of correct retrieval from 982 test set images.
+C.2. Retrieval Heatmap
+We visualized the retrieval heatmap for Subject 1 on all 982 test images of the NSD dataset in Figure 9.
+It can be observed that the similarity is highest on the diagonal, and the color of the retrieved heat map is darker.
+It shows that Lite-Mind has effectively retrieved corresponding images, even if there are many similar images in the test set, which verifies the fine-grained ability of Lite-Mind.
+C.3. Information Alignment
+We visualized the T-SNE plot between the voxel embeddings output by DFT Backbone and the image embeddings of frozen CLIP as the accuracy of the test set improved, as shown in Figure 13.
+We can observe that as the training progresses, the retrieval accuracy of the test set improves, and the shape of voxel embeddings tends to be closer to image embeddings, indicating the success of contrastive learning.
+C.4. More Retrieval Cases
+In this section, we visualize retrieval failure cases in Figure 10, although only dozens of images are not in the Top 1.
+From the left half of the Figure, it can be seen that though not in Top 1, Lite-Mind still retrieved ground-truth in Top 2, and images in Top 4 are similar(either semantically similar or structurally similar, eg.
+animals in the wild or an airplane on the runway).
+On the contrary, a smaller proportion of ground-truth did not appear within the Top 4, as shown in the right half of the Figure.
+From the perspective of the images themselves, most of the scenes are too complex, even abstract (as shown in the second image), which may be the reason for the retrieval failure.
+Figure 10. Partial failure retrieval results of Lite-Mind on all 982 test images for Subject 1. The number below each image represents the similarity score.
+C.5. Visualization in Frequency Domain
+We visualized the weights of Filter library of Filter Blocks, as shown in Figure 11 and 12.
+Visualization is divided into the real part (left) and imaginary part (right) of filter weights.
+It can be observed that different filters have varying degrees of attention to different tokens, and the frequency domain better captures this characteristic.
+Interestingly, the weight of the imaginary part for the first and last tokens is almost always 0, indicating that the noise is distributed in these two tokens.
+(a) The real part and imaginary of a filter weights.
+(b) The real part and imaginary of another filter weights from the same Filter Library.
+Figure 11. Weights visualization for DFT Backbone of $257\times 768$ embedding length.
+(a) The real part and imaginary of a filter weights.
+(b) The real part and imaginary of another filter weights from the same Filter Library.
+Figure 12. Weights visualization for DFT Backbone of 768 CLS embedding length.
+Appendix D Theoretical Analysis.
+D.1. Complex Multiplication.
+For two complex number values $\mathcal{Z}_{1}=(a+jb)$ and $\mathcal{Z}_{2}=(c+jd)$, where $a$ and $c$ is the real part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively, $b$ and $d$ is the imaginary part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively.
+Then the multiplication of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ is calculated by:
+(10) $\mathcal{Z}_{1}\mathcal{Z}_{2}$ $=(a+jb)(c+jd)$
+$=ac+j^{2}bd+jad+jbc$
+$=(ac-bd)+j(ad+bc)$
+D.2. Theorem Proof.
+Theorem 1.
+Suppose that H is the representation of raw fMRI voxel tokens and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of voxel tokens in the spatial domain is equal to the energy of its representation in the frequency domain.
+Formally, we can express this with the above notations:
+(11) $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$
+Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the token dimension, $f$ is the frequency dimension.
+Proof.
+Given the representation of raw voxel token series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
+(12) $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{\textbf{H}(v)}{\textbf{H}^{*}(v)}\,\mathrm{d}v$
+where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
+(13) $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v$ $=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v$
+$=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f$
+$=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f$
+$=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$
+Proved. ∎
+Therefore, the energy of a voxel token series in the spatial domain is equal to the energy of its representation in the frequency domain.
+Theorem 2.
+Given the voxel token series input H and its corresponding frequency domain conversion H, the operations of frequency-domain MLP on H can be represented as global convolutions on H in the spatial domain.
+This can be given by:
+(14) $\mathcal{H}\mathcal{W}+\mathcal{B}=\mathcal{F}(\textbf{H}*W+B)$
+Where $\mathcal{F}$ is DFT, * is a circular convolution, $\mathcal{W}$ and $\mathcal{B}$ are the complex number weight and bias, and $W$ and $B$ are the weight and bias in the spatial domain.
+Proof.
+Suppose that we conduct operations in the N (i.e., token dimension), then
+(15) $\mathcal{F}(\textbf{H}(v)*W(v))=\int_{-\infty}^{\infty}{(\textbf{H}(v)*W(v))e^{-j2\pi fv}}\,\mathrm{d}v$
+According to convolution theorem, $\textbf{H}(v)*W(v)=\int_{-\infty}^{\infty}(\textbf{H}(\tau)W(v-\tau))\,\mathrm{d}\tau$, then
+(16) $\mathcal{F}(\textbf{H}(v)*W(v))$ $=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{(\textbf{H}(\tau)W(v-\tau))e^{-j2\pi fv}}\,\mathrm{d}\tau}\,\mathrm{d}v$
+$=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(v-\tau)e^{-j2\pi fv}}\,\mathrm{d}v}\textbf{H}(\tau)\,\mathrm{d}\tau$
+Let $x=v-\tau$,then
+(17) $\mathcal{F}(\textbf{H}(v)*W(v))$ $=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi f(x+\tau)}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau$
+$=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi fx}e^{-j2\pi f\tau}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau$
+$=\int_{-\infty}^{\infty}{(\textbf{H}(\tau)*)e^{-j2\pi f\tau}}\,\mathrm{d}\tau\int_{-\infty}^{\infty}{(\textbf{W}(x)*)e^{-j2\pi fx}}\,\mathrm{d}x$
+$=\mathcal{H}(f)\mathcal{W}(f)$
+Accordingly, $\textbf{H}(v)*W(v)$ in the spatial domain is equal to $\mathcal{H}(f)\mathcal{W}(f)$ in the frequency domain.
+Therefore, the operations of FreMLP ($\mathcal{H}\mathcal{W}+\mathcal{B}$) in the token dimension (i.e., v = N ) are equal to the operations ($\mathcal{F}(\textbf{H}*W+B)$) in the spatial domain.
+This implies that frequency-domain MLPs can be viewed as global convolutions in the spatial domain.
+Proved.
+∎
+D.3. Complexity Analysis
+For a fMRI voxel with a length of $l$, we divide it into $n$ patches.
+Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP Backbone and DFT Backbone respectively, the middle layer dimension of MLP Backbone is $D$, and the alignment embedding dimension is $n^{\prime}\times D^{\prime}$, where $n^{\prime}$ is the the number of tokens of CLIP.
+The time complexity of MLP Backbone is $O(lD+L_{1}D^{2}+n^{\prime}DD^{\prime})$.
+For DFT Backobone, the time complexity of patchify and tokenization is $O(lD^{\prime})$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nD^{\prime}\operatorname{log}n+nD^{\prime})$.
+The time complexity of FreMLP is $O(2nD^{\prime}\operatorname{log}n+2nn^{\prime}D^{\prime}+2n^{\prime}D^{\prime})$.
+Thus the time complexity of the entire DFT Backbone is:
+(18) $O(4nD^{\prime}\operatorname{log}n+(n+2nn^{\prime}+2n^{\prime})D^{\prime})$
+$=O((n\operatorname{log}n+nn^{\prime}+n^{\prime})D^{\prime})$
+Quantitative analysis algorithm complexity for DFT Backbone has been shown in Table 3.
+Figure 13.
+T-SNE visualization between the voxel embeddings output by DFT Backbone and the image embedding of frozen CLIP.
+Accuracy in the title refers to the hit rate of correct retrieval from 982 test set images and the blue dots represent voxel embeddings.